### v2

Abstract
Recent advances in learned image compression (LIC) have achieved remarkable performance improvements over traditional codecs. Notably, the MLIC series—LICs equipped with multi-reference entropy models—have substantially surpassed conventional image codecs such as Versatile Video Coding (VVC) Intra. However, existing MLIC variants suffer from several limitations: performance degradation at high bitrates due to insufficient transform capacity, suboptimal entropy modeling that fails to capture global correlations in initial slices, and lack of adaptive channel importance modeling. In this paper, we propose MLICv2 and MLICv2+, enhanced successors that systematically address these limitations through improved transform design, advanced entropy modeling, and exploration of the potential of instance-specific optimization. For transform enhancement, we introduce a lightweight token mixing block inspired by the MetaFormer architecture, which effectively mitigates high-bitrate performance degradation while maintaining computational efficiency. For entropy modeling improvements, we propose hyperprior-guided global correlation prediction to extract global context even in the initial slice of latent representation, complemented by a channel reweighting module that dynamically emphasizes informative channels. We further explore enhanced positional embedding and guided selective compression strategies for superior context modeling. Additionally, we apply the Stochastic Gumbel Annealing (SGA) to demonstrate the potential for further performance improvements through input-specific optimization. Extensive experiments demonstrate that MLICv2 and MLICv2+ achieve state-of-the-art results, reducing Bjøntegaard-Delta Rate by 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% on Kodak, Tecnick, and CLIC Pro Val datasets, respectively, compared to VTM-17.0 Intra.
1 Introduction
Image compression constitutes a fundamental challenge in multimedia communication, with the objective of achieving compact representation while preserving reconstruction quality. Traditional compression methods, exemplified by VVC Intra [12], rely on block-based hybrid coding frameworks refined over decades. However, as conventional approaches near performance saturation, learned image compression (LIC) has emerged as a transformative paradigm, demonstrating substantial potential to surpass hand-crafted codecs through end-to-end optimization.
Contemporary LIC methods [9, 64, 10, 50, 51, 16, 75, 31, 30, 88, 35, 86, 68, 87, 38, 19, 71, 26, 74, 78, 53, 15, 45, 34, 33, 32] predominantly adopt autoencoder-based frameworks with rate-distortion optimization, comprising analysis transforms, quantization, synthesis transforms, and entropy models. The analysis transform maps input images to latent representations, which undergo quantization before reconstruction via synthesis transforms. Entropy models estimate quantized latent distributions to facilitate efficient entropy coding. Advanced techniques of transform architectures and entropy models have enabled several LICs [30, 34, 33, 32, 45, 40, 47, 20] to achieve performance comparable to or exceeding advanced traditional codecs like VVC Intra [12].
Figure 1: BD-Rate and decoding time during inference on CLIC Professional Valid [65] with 2K resolution. The bubble size indicates the number of model parameters.
Entropy modeling plays a pivotal role in LIC performance, aiming to minimize conditional entropy through informative prior capture. While hyperprior models [10] provide side information, context models [52, 31] capture spatial dependencies between previously decoded and current elements. Building on these foundations, Jiang et al. introduced the MLIC series [34, 33], which leverages multi-reference priors through latent representation partitioning into channel-wise slices [51], capturing both inter-slice and intra-slice dependencies. MLIC employs checkerboard partitioning [31] within slices for two-pass decoding, while attention mechanisms model global dependencies from previous slices. MLIC++ [33] addressed computational complexity by replacing vanilla attention [70] with linear attention [60] and modeling global correlations across previous slices.
Despite the strong overall performance of the MLIC series, several critical limitations persist and motivate our work. First, regarding transform design, MLIC models exhibit performance degradation at high bitrates, primarily attributed to insufficient transform capacity for preserving latent representation information. Second, concerning entropy modeling, global correlation estimation relies exclusively on previous slices, precluding global correlation capture in initial slices. Moreover, existing approaches treat all channels equally without considering varying informational importance, potentially limiting entropy modeling accuracy.
To systematically address these limitations, we propose MLICv2, featuring enhanced transform and entropy modeling modules, and MLICv2+, which additionally incorporates iterative refinement techniques [80] to demonstrate framework potential.
For transform improvement, we design an efficient token mixing block replacing residual blocks in MLIC++. Inspired by MetaFormer architecture [81], our block employs depth-wise convolution for spatial token mixing and lightweight gated interaction for point-wise operations, achieving superior high-bitrate performance with reduced complexity.
For entropy modeling enhancement, we propose MEMv2, an advanced multi-reference entropy model capturing diverse dependencies. We introduce hyperprior-guided global correlation prediction for initial slices, leveraging spatial distribution similarity between hyperprior and latent representations. Our two-stage context modeling strategy performs spatial aggregation followed by adaptive channel-wise reweighting to capture feature importance variations. Additional improvements include 2D Rotary Positional Embedding (RoPE) and guided selective compression for enhanced modeling and coding efficiency.
Additionally, we employ Stochastic Gumbel Annealing (SGA) [80] to iteratively refine latent representations and side information for each input. This approach demonstrates our framework’s potential for input-specific rate-distortion optimization while preserving decoder complexity, suggesting opportunities for further architectural enhancements.
Our comprehensive approach yields significant improvements, with MLICv2 and MLICv2+ substantially outperforming MLIC++ as demonstrated in Figure 1. MLICv2 achieves BD-rate reductions of 16.54%, 21.61%, and 16.05% on Kodak, Tecnick, and CLIC Pro Val datasets respectively, while MLICv2+ further improves these results to 20.46%, 24.35%, and 19.14% over VTM-17.0 Intra. Our contributions are summarized as follows:
  • We introduce MetaFormer-based architecture for LIC through our simple token mixing block, achieving superior high-bitrate performance with reduced computational complexity compared to previous residual blocks.
  • We systematically improve entropy modeling through multiple improvements: (i) hyperprior-guided global correlation prediction addressing initial slice limitations, (ii) the first two-stage context modeling strategy with spatial aggregation and adaptive channel reweighting, and (iii) advanced positional embedding and selective compression mechanisms.
  • We demonstrate the effectiveness of our framework by applying iterative refinement techniques (SGA), showing consistent performance gains without decoder complexity increase and suggesting opportunities for further architectural enhancements.
  • Extensive experiments validate our approach, with MLICv2 and MLICv2+ achieving BD-rate reductions of 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% over VTM-17.0 Intra on Kodak, Tecnick, and CLIC Pro Val datasets respectively.
2 Related Works
2.1 Learned Image Compression
Autoencoder-based learned image compression [9, 64, 50, 16, 30, 34, 33, 51, 32] follows the transform coding paradigm, consisting of an analysis transform $g_{a}$ with parameters $\phi$, a synthesis transform $g_{s}$ with parameters $\psi$, a quantization function $\lceil\cdot\rfloor$, and an entropy model $p$ with parameters $\rho$. Given an input image $\boldsymbol{x}\sim p_{\boldsymbol{x}}$, the process is formulated as:
$\boldsymbol{y}=g_{a}(\boldsymbol{x};\phi),\hat{\boldsymbol{y}}=\lceil\boldsymbol{y}\rfloor,\hat{\boldsymbol{x}}=g_{s}(\hat{\boldsymbol{y}};\psi).$ (1)
The rate-distortion optimization is achieved through:
$\mathcal{L}=\mathbb{E}_{\boldsymbol{x}\sim p_{\boldsymbol{x}}}\left[\underbrace{-\log_{2}\int p(\hat{\boldsymbol{y}};\rho)d\hat{\boldsymbol{y}}}_{\textrm{bit-rate}}+\lambda\times\underbrace{\mathcal{D}(\hat{\boldsymbol{x}},\boldsymbol{x})}_{\textrm{distortion}}\right],$ (2)
where $\lambda$ balances the rate-distortion trade-off. During training, quantization is approximated by adding uniform noise [9] or straight-through estimation [64]. Ballé et al. [10] introduce hyperprior networks using hyper analysis $h_{a}$ and hyper synthesis $h_{s}$ to improve entropy estimation through quantized side information $\hat{\boldsymbol{z}}$ extracted from latent representations.
Transform architectures have evolved significantly, incorporating Generalized Divisive Normalization (GDN) [8], residual non-local attention [85, 15], deep residual networks [16], and transformer-based designs [87, 88, 48]. Mixed CNN-transformer [45] and Mamba [82, 25, 56] architectures have also been explored for capturing both local and non-local interactions. Entropy estimation has progressed from simple univariate Gaussian models [10] to sophisticated approaches including mean-scale Gaussian models [50], asymmetric Gaussian distributions [17], Gaussian mixture models [16], and generalized Gaussian models [83], significantly improving compression efficiency.
2.2 Conditional Entropy Models
Conditional entropy models leverage context $\hat{\boldsymbol{C}}$ to reduce bit-rate, exploiting the principle that conditional entropy is bounded by unconditional entropy:
[…]
Local context has evolved from serial approaches like PixelCNN [69, 50] to efficient parallel methods including zigzag scanning [41] and checkerboard partitioning [30]. These intra-slice local contexts are complemented by inter-slice approaches and channel-wise autoregressive coding [51] for faster decoding.
Recent works investigate long-range correlations among distant elements [55, 26, 36, 34, 33]. Qian et al. [55, 54] employ neighbor elements and bi-directional attention for global correlation estimation. Guo et al. [26] utilize L-2 distance metrics, while Kim et al. [36] transmit global information to decoders. Recently, Jiang et al. [34] propose MLIC and MLIC+, dividing global contexts into inter-slice and intra-slice components. The latent representation is partitioned into multiple slices for autoregressive conditional coding, employing checkerboard attention for intra-slice contexts and convolutional layers for inter-slice contexts. However, the quadratic complexity of vanilla attention [70] limits scalability for high-resolution images, which is addressed by MLIC++ [33] through linear attention mechanisms [60].
Figure 2: The overall architecture of MLICv2/MLICv2+. $g_{a}$ is the analysis transform, $g_{s}$ is the synthesis transform, $h_{a}$ is the hyper analysis, and $h_{s}$ is the hyper synthesis. $g_{ep}$ is the entropy parameter module. $f_{s}$ is the selective compression predictive module. $g_{lrp}$ is the latent residual prediction module [51]. AE and AD are arithmetic encoding and decoding. $\uparrow$ denotes upsampling and $\downarrow$ denotes downsampling. $\boldsymbol{x}$} is the input image and $\hat{\boldsymbol{x}}$ is the reconstructed image. $\boldsymbol{y}$ is the latent representation and $\hat{\boldsymbol{y}}$ is the quantized latent representation. $\hat{\boldsymbol{y}}^{i}$ is the $i$-th slice of $\hat{\boldsymbol{y}}$. $\boldsymbol{\mu}^{i},\boldsymbol{\sigma}^{i}$ are the mean and scale of $\hat{\boldsymbol{y}}^{i}$. $\boldsymbol{s}^{i}$ is the predicted selective coding map. $\boldsymbol{r}^{i}$ is the predicted residual. $\hat{\boldsymbol{z}}$ is the side information. $M,N$ are the channel numbers, which are $320,192$ in MLICv2/MLICv2+. “Refine” is employed in MLICv2+.
3 Method
3.1 Background: MLIC++
Our proposed MLICv2 and MLICv2+ build upon the foundation established by MLIC++ [33]. MLIC++ employs a simplified version of Cheng’20’s transform [16], where attention modules are removed to reduce computational complexity. For entropy modeling, MLIC++ introduces a linear-complexity multi-reference entropy model (MEM++), which effectively captures hyperprior, intra-slice, and inter-slice dependencies.
The hyperprior information $\hat{\boldsymbol{H}}$ is derived from side information $\hat{\boldsymbol{z}}$. Following the approach of Minnen et al. [51], the latent representation $\hat{\boldsymbol{y}}$ is partitioned into $L+1$ slices: $\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\ldots,\hat{\boldsymbol{y}}^{L}$}. Each slice $\hat{\boldsymbol{y}}^{i}$ is further divided into anchor $\hat{\boldsymbol{y}}^{i}_{ac}$ and non-anchor $\hat{\boldsymbol{y}}^{i}_{na}$ components using checkerboard partitioning [31].
The inter-slice local context module $g_{\ell}^{inter}$ captures inter-slice local context $\hat{\boldsymbol{C}}^{<i}_{\ell}$ from preceding slices $\hat{\boldsymbol{y}}^{<i}$. To model inter-slice global context $\hat{\boldsymbol{C}}^{<i}_{g}$, MLIC++ employs linear attention [60] to achieve an optimal balance between performance and computational complexity. The intra-slice local context $\hat{\boldsymbol{C}}^{i}_{\ell}$ is extracted from the anchor partition $\hat{\boldsymbol{y}}^{i}_{ac}$ using overlapped window-based checkerboard attention $g_{\ell}^{intra}$.
A key insight in MLIC++ is that feature maps from different slices exhibit similar global similarity patterns, as they can be viewed as thumbnails of the same content. Therefore, the attention maps between anchor and non-anchor parts of the (i1)(i-1)-th slice are leveraged to predict global similarities for the current slice. The intra-slice global context $\hat{\boldsymbol{C}}_{g}^{i}$ is extracted from $\hat{\boldsymbol{y}}^{i}_{ac}$ via the predicted attention map and the intra-slice global context module $g_{g}^{intra}$. This process is formulated as:
$\hat{\boldsymbol{C}}^{\textless i}_{\ell}$ $=g_{\ell}^{inter}(\hat{\boldsymbol{y}}^{\textless i}),\hat{\boldsymbol{C}}^{\textless i}_{g}=g_{g}^{inter}(\hat{\boldsymbol{y}}^{\textless i}),$ (4)
$\hat{\boldsymbol{C}}^{i}_{\ell}$ $=g_{\ell}^{intra}(\hat{\boldsymbol{y}}^{i}_{ac}),\hat{\boldsymbol{C}}^{i}_{g}=g_{g}^{intra}(\hat{\boldsymbol{y}}^{i-1}_{ac},\hat{\boldsymbol{y}}^{i-1}_{na},\hat{\boldsymbol{y}}^{i}_{ac}).$
The estimated means $\boldsymbol{\mu}^{i}_{ac}$, $\boldsymbol{\mu}^{i}_{na}$ and scales $\boldsymbol{\sigma}^{i}_{ac}$, $\boldsymbol{\sigma}^{i}_{na}$ are extracted from the contextual information using the entropy parameter module $g_{ep}$. MLIC++ incorporates mixed quantization [51] and the latent residual prediction module $g_{lrp}$ [51]. In mixed quantization, zero-centered straight-through estimation (STE) [64] is applied for distortion computation, while additive uniform noise (AUN) [9] is used for entropy estimation during training. The module $g_{lrp}$ predicts quantization residuals based on decoded slices and hyperprior information:
$\hat{\boldsymbol{y}}_{ac}^{i}$ $=\textrm{STE}(\boldsymbol{y}^{i}_{ac}-\boldsymbol{\mu}^{i}_{ac})+\boldsymbol{\mu}^{i}_{ac}+\boldsymbol{r}^{i}_{ac},$ (5)
$\hat{\boldsymbol{y}}_{na}^{i}$ $=\textrm{STE}(\boldsymbol{y}^{i}_{na}-\boldsymbol{\mu}^{i}_{na})+\boldsymbol{\mu}^{i}_{na}+\boldsymbol{r}^{i}_{na},$
where
[…]
3.2 Overall Architecture
Figure 2 illustrates the overall architecture of our proposed MLICv2 and MLICv2+. MLICv2 preserves the fundamental structure of MLIC++ while introducing several key enhancements to improve compression efficiency, particularly at high bit rates. First, we replace residual blocks following downsampling and upsampling operations with lightweight token mixing blocks, which enhance representational capacity while reducing computational complexity. Second, we upgrade the multi-reference entropy model to MEMv2 by incorporating enhanced inter-slice and intra-slice context modules with context reweighting and advanced positional embedding to better capture latent dependencies. Third, we introduce a selective compression module $f_{s}$ that intelligently predicts and skips zero elements in latent representations, reducing both bit rate and arithmetic coding time without compromising reconstruction quality. To demonstrate the potential for input-specific optimization, we propose MLICv2+, which iteratively refines latent features $\boldsymbol{y}$ and side information $\boldsymbol{z}$ based on the rate-distortion characteristics of individual inputs, enabling tailored compression strategies for different content types.
Figure 3: Visualization of the channels of hyperprior $\hat{\boldsymbol{H}}$ ($\lambda=0.013$). $\hat{\boldsymbol{H}}$ extracted from the latent representation of Image Kodim20 exhibits similar global similarity to original Kodim20.
3.3 Efficient Simple Token Mixing Transform
As illustrated in Figure 6, MLIC++ exhibits performance degradation at high bit rates due to limited capacity of transform modules to retain sufficient information in latent representations [76]. The representational capacity is fundamentally influenced by the number of channels. As demonstrated by Ballé et al. [6], when channel numbers are limited, the performance gap between low- and high-capacity models is negligible at low bit rates. However, this gap becomes pronounced as bit rates increase, with higher-capacity models achieving significantly better reconstruction quality.
Simply increasing channel numbers often results in prohibitive computational complexity. A more practical approach is to enhance the architectural design itself to improve transform capacity without substantially increasing computational cost. Recent work has adopted transformer layers [46] for transform coding [45, 40, 88, 48], demonstrating remarkable performance improvements. However, existing techniques [45] can introduce up to twice the complexity of original residual blocks in the MLIC series.
Inspired by the elegance of MetaFormer [81], which consists of token mixing and channel-wise interaction components, we investigate its effectiveness for LIC. However, MetaFormer’s use of pooling for token mixing, while efficient, is suboptimal for compression tasks due to information loss. We propose a simple token mixing block as illustrated in Figure 2. For token mixing, we employ a depth-wise residual block (DepthRB) [32] for nonlinear embedding, depth-wise convolution for spatial token mixing, and a gate block [32] for efficient channel-wise interactions. LayerNorm (LN) [5] is adopted for stable training. Given input feature $\boldsymbol{\omega}$, the process is formulated as:
$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\textrm{Conv}1\times 1(\textrm{DWConv}5\times 5(\textrm{DepthRB}(\textrm{LN}(\boldsymbol{\omega})))),$ (7)
$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\textrm{Gate}(\textrm{LN}(\boldsymbol{\omega})),$
where $\boldsymbol{\omega}\in\mathbb{R}^{c\times h\times w}$ represents the input feature with $c$, $h$, $w$ denoting channels, height, and width respectively. Stacking two token mixing blocks creates deeper transforms with enhanced expressiveness. Compared to previous residual blocks [16] consisting of two $3\times 3$ convolutions, our token mixing block achieves lower complexity while delivering significantly better performance, especially at high bit rates. The complexity analysis reveals that stacking two simple token-mixing blocks reduces multiply-accumulate operations (MACs) by 7.38% at $c=192$ and by 8.87% at $c=320$, compared with stacking a single residual block [16, 34, 33].
3.4 Enhanced Multi-Reference Entropy Modeling
The enhanced multi-reference entropy model (MEMv2) is presented in Figure 2. To address the limitations of MLIC++, we introduce several key innovations: hyperprior-based global correlation prediction, context reweighting to handle varying channel importance, 2D Rotary Positional Embedding (RoPE) and guided selective compression for improved context modeling and entropy coding.
3.4.1 Hyperprior-Guided Global Correlation Prediction
In MLIC++ [33], global similarity between anchor and non-anchor parts of the previous slice is exploited to predict current slice similarities. However, for the first slice $\hat{\boldsymbol{y}}^{0}$, no such reference exists. The only available conditional information is the hyperprior $\hat{\boldsymbol{H}}$, obtained by upsampling side information $\hat{\boldsymbol{z}}$ extracted from latent representation $\boldsymbol{y}$. As illustrated in Figure 3, the hyperprior preserves global similarity patterns similar to the original image. We leverage this property by estimating global similarity between $\hat{\boldsymbol{y}}_{ac}^{0}$ and $\hat{\boldsymbol{y}}_{na}^{0}$ using the similarity between anchor and non-anchor parts of the hyperprior, $\hat{\boldsymbol{H}}_{ac}$ and $\hat{\boldsymbol{H}}_{na}$. Following MLIC++, we employ linear attention [60] for global context modeling and introduce a Gate block [32] for enhanced pointwise interaction:
$\hat{\boldsymbol{C}}_{g}^{0}$ $=\textrm{Softmax}_{2}(\hat{\boldsymbol{H}}_{na,q})\left(\textrm{Softmax}_{1}(\hat{\boldsymbol{H}}_{ac,k})^{\top}\hat{\boldsymbol{y}}^{0}_{g,v}\right),$ (8)
$\hat{\boldsymbol{C}}_{g}^{0}$ $=\textrm{Gate}(\hat{\boldsymbol{C}}_{g}^{0})+\hat{\boldsymbol{C}}_{g}^{0},$
where $\hat{\boldsymbol{H}}_{na,q}=\text{Linear}(\hat{\boldsymbol{H}}_{na})$, $\hat{\boldsymbol{H}}_{ac,k}=\text{Linear}(\hat{\boldsymbol{H}}_{ac})$, and $\hat{\boldsymbol{y}}^{0}_{g,v}=\text{Linear}(\hat{\boldsymbol{y}}_{ac}^{0})$.
While prior methods [30, 34, 33, 32] capture only local dependencies in the first slice, our hyperprior-guided mechanism models global dependencies effectively. Given the information compaction property of latent representations [30], the fact that the first slice typically exhibits higher entropy as it serves as reference for subsequent slices, incorporating global context significantly reduces bit rate and improves compression performance.
Figure 4: Proposed context reweighting module.
3.4.2 Context Reweighting
In MLIC++ [33], the contexts are extracted by spatial attention mechanism. Given the query $\boldsymbol{Q}\in\mathbb{R}^{L\times c}$, key $\boldsymbol{K}\in\mathbb{R}^{L\times c}$ and the value $\boldsymbol{V}\in\mathbb{R}^{L\times c}$, where $L=hw$ is the sequence length, $h$ is the height, $w$ is the width, and $c$ is the channel number, in most spatial attention [70, 60], the attention map is computed by $\textrm{Sim}(\boldsymbol{Q}\boldsymbol{K}^{\top})\in\mathbb{R}^{L\times L}$, where Sim is the predefined function, such as Softmax [70] or two independent Softmax operations [60], indicating the information is gathered and weighted from spatial tokens ($\in\mathbb{R}^{1\times c}$). However, one common limitation is that all channel-wise features are treated equally, without considering that different channels may carry varying levels of importance. We argue that adaptively reweighting channel-wise context features can lead to more accurate entropy modeling. By suppressing noisy or less informative channels and emphasizing more relevant ones, the model has the potential to capture context information more precisely and improve overall compression performance, which has not been considered in existing methods [34, 33, 30, 31, 51, 19, 22, 54, 55].
To address this issue, we introduce a two-stage context modeling strategy, where spatial aggregation [33] is first performed, followed by adaptive channel-wise context reweighting to capture the varying importance of different features. The architecture is in Fig. 4. Specifically, in context reweighting, the context is transformed to query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$, and the query and values are transposed to compute the channel-wise attention map [3]. The process is formulated as:
$\boldsymbol{O}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})\boldsymbol{V}^{\top}.$
The resolution of the attention map $\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})$ is $c\times c$, indicating the importance of each channel ($\in\mathbb{R}^{L\times 1}$). The complexity of Eq. (8) is $\mathcal{O}(c^{2}hw)$. $\boldsymbol{O}$ is then fed into a Gate block [32] for efficient point-wise interactions. This attention-based reweighting helps the model focus on the most relevant parts of the context. To the best of our knowledge, this is the first attempt to explicitly model channel-wise importance in context features for LIC. Our approach brings consistent performance improvements with minimal additional cost.
3.4.3 2D Rotary Position Embedding
MLIC++ [33] applies relative position embedding (RPE) from Swin-Transformer [46] directly to attention maps as additional bias. However, this direct addition may limit interaction with attention weights. Recently, Rotary Positional Embedding (RoPE) [63], a multiplication-based approach, has been widely adopted in language models [66, 58]. RoPE offers several advantages: compatibility with linear attention (whereas RPE requires adding biases that conflict with implicit attention computation), better extrapolation capabilities, and more concise implementation. For 2D image content, we extend the original 1D RoPE to handle horizontal, vertical, and diagonal relationships. Given query $\boldsymbol{q}_{m_{x},m_{y}}$ and key $\boldsymbol{k}_{n_{x},n_{y}}$ at 2D positions $(m_{x},m_{y})$ and $(n_{x},n_{y})$, we rotate queries and keys twice with different angles $\theta_{x}$ and $\theta_{y}$ for horizontal and vertical directions. When $c=2$, the rotation matrix for the query becomes:
$\underbrace{\begin{bmatrix}\cos m_{x}\theta_{x}&-\sin m_{x}\theta_{x}\\ \sin m_{x}\theta_{x}&\cos m_{x}\theta_{x}\end{bmatrix}}_{\textrm{horizontal}}\underbrace{\begin{bmatrix}\cos m_{y}\theta_{y}&-\sin m_{y}\theta_{y}\\ \sin m_{y}\theta_{y}&\cos m_{y}\theta_{y}\end{bmatrix}}_{\textrm{vertical}}=\begin{bmatrix}\cos(m_{x}\theta_{x}+m_{y}\theta_{y})&-\sin(m_{x}\theta_{x}+m_{y}\theta_{y})\\ \sin(m_{x}\theta_{x}+m_{y}\theta_{y})&\cos(m_{x}\theta_{x}+m_{y}\theta_{y})\end{bmatrix}$. (9)
The attention score computation yields:
$\boldsymbol{q}_{m_{x},m_{y}}^{\prime}=\boldsymbol{q}_{m_{x},m_{y}}\begin{bmatrix}\cos(m_{x}\theta_{x}+m_{y}\theta_{y})&-\sin(m_{x}\theta_{x}+m_{y}\theta_{y})\\ \sin(m_{x}\theta_{x}+m_{y}\theta_{y})&\cos(m_{x}\theta_{x}+m_{y}\theta_{y})\end{bmatrix}$, (10)
$\boldsymbol{k}_{n_{x},n_{y}}^{\prime}=\boldsymbol{k}_{n_{x},n_{y}}\begin{bmatrix}\cos(n_{x}\theta_{x}+n_{y}\theta_{y})&-\sin(n_{x}\theta_{x}+n_{y}\theta_{y})\\ \sin(n_{x}\theta_{x}+n_{y}\theta_{y})&\cos(n_{x}\theta_{x}+n_{y}\theta_{y})\end{bmatrix}$,
$\boldsymbol{q}_{m_{x},m_{y}}^{\prime}(\boldsymbol{k}_{n_{x},n_{y}}^{\prime})^{\top}=\boldsymbol{q}_{m}\begin{bmatrix}\cos\theta_{(m_{x}-n_{x},m_{y}-n_{y})}&-\sin\theta_{(m_{x}-n_{x},m_{y}-n_{y})}\\ \sin\theta_{(m_{x}-n_{x},m_{y}-n_{y})}&\cos\theta_{(m_{x}-n_{x},m_{y}-n_{y})}\end{bmatrix}\boldsymbol{k}_{n}^{\top},$
where $\theta_{(m_{x}-n_{x},m_{y}-n_{y})}=(m_{x}-n_{x})\theta_{x}+(m_{y}-n_{y})\theta_{y}$ encodes relative positions. Due to the ambiguity in defining relative importance of horizontal versus vertical positional dependencies, we make angles $\theta_{x}$ and $\theta_{y}$ learnable, initializing them to 10000 following original RoPE [63].
3.4.4 Guided Selective Compression
Zero-centered quantization in Equation 5 produces many zero elements in $\lceil\boldsymbol{\Lambda}\rfloor=\lceil\boldsymbol{y}-\boldsymbol{\mu}\rfloor$, as illustrated in Figure 5. While these zeros consume minimal bit rate, they account for significant arithmetic encoding/decoding time.
Recent selective compression methods [39, 2] have employed neural networks to predict selective compression maps, but they are primarily applied to simpler entropy models. In addition, scale-based prediction has been introduced in recent works [62, 84]. However, when extended to advanced entropy models [34, 33], the training phase becomes highly unstable, with the overall loss value becoming “nan” after several steps, regardless of whether the parameters of the compression model are frozen or not. The observed training instability is primarily caused by the dependency introduced by the context modeling, rather than by the selective compression mechanism itself. Specifically, skipping non-zero elements in the current slice can significantly affect the entropy modeling of subsequent slices due to the inter-latent dependencies established by the context model. Considering the substantial performance gains enabled by context modeling, our goal is to skip the encoding of zero elements as much as possible—on top of the existing context modeling framework [51, 30, 33, 34]—in order to reduce the time spent on arithmetic encoding and decoding.
To address above issues, the guided selective compression module $f_{s}$ is proposed for post-training. This module is employed on the converged MLICv2, with the weights of MLICv2 frozen for stable training. The architecture of the guided selective compression $f_{s}$ module is presented in Fig. 2. Specifically, in $f_{s}$, the scale-based selective prediction is employed as the initial predicted selective compression map. For $j$-th element ${s}_{\boldsymbol{\sigma},j}$ in $\boldsymbol{s}_{\boldsymbol{\sigma}}$, the scale-based selective compression value is:
${s}_{\boldsymbol{\sigma},j}=\left\{\begin{aligned} &0,\quad{\sigma}_{j}\textless{\xi},\\ &1,\quad{\sigma}_{j}\geq{\xi},\end{aligned}\right$. (11)
where $\xi$ is a predefined threshold, set to $0.3$. For $\boldsymbol{y}^{i}_{ac}$, the decoded elements of previous slices serve as priors. For $\boldsymbol{y}^{i}_{na}$, the decoded elements of previous slices and $\boldsymbol{y}^{i}_{ac}$ serve as priors. The hyperprior $\hat{\boldsymbol{H}}$ also serves as a prior. The predicted selective map is formulated as:
$\boldsymbol{\varepsilon}^{i}_{ac}$ $=f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},ac}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\hat{\boldsymbol{H}}\right),\boldsymbol{s}^{i}_{ac}=\textrm{STE}\left(\textrm{Sigmoid}\left(\boldsymbol{\varepsilon}^{i}_{ac}\right)\right),$ (12)
$\boldsymbol{\varepsilon}^{i}_{na}$ $=f_{s}\Bigl{(}\boldsymbol{s}_{\boldsymbol{\sigma},na}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\boldsymbol{s}^{i}_{ac},\hat{\boldsymbol{H}}\Bigr{)},\boldsymbol{s}^{i}_{na}=\textrm{STE}\left(\textrm{Sigmoid}\left(\boldsymbol{\varepsilon}^{i}_{na}\right)\right).$
During training, $\lceil\boldsymbol{\Lambda}\rfloor$ is known, therefore cross-entropy loss is employed to optimize $f_{s}$, treating selective prediction as a binary classification task. Additionally, elements with larger values have a higher impact on distortion, so the cross-entropy loss is weighted by $\lceil\boldsymbol{\Lambda}\rfloor+1$. To the best of our knowledge, we are the first to highlight the importance of quantization residuals and incorporate a weighted loss accordingly. The overall loss is:
Figure 5: Ratios of different elements in $\lceil\boldsymbol{\Lambda}\rfloor=\lceil\boldsymbol{y}-\boldsymbol{\mu}\rfloor$ when the model is optimized for MSE and the $\lambda=0.013$. There are many zero elements in $\lceil\boldsymbol{\Lambda}\rfloor$.
$\mathcal{L}_{s}=\sum_{i=0}^{L}\Bigl{(}(\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor+1)\odot\mathcal{L}^{i}_{ce,ac}+(\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor+1)\odot\mathcal{L}^{i}_{ce,na}\Bigr{)},$ (13)
$\mathcal{L}_{ce,ac}^{i}=-\frac{\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\log\boldsymbol{\varepsilon}^{i}_{ac}+\left(1-\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\right)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{ac})}{2},$
$\mathcal{L}_{ce,na}^{i}=-\frac{\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor\odot\log\boldsymbol{\varepsilon}^{i}_{na}+\left(1-\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor\right)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{na})}{2}.$
3.5 Iterative Latent Refinement
To validate our MLICv2’s potential and demonstrate achievable performance improvements, we employs encoding-time scaling to find optimal quantized latent representation $\hat{\boldsymbol{y}}$ and side information $\hat{\boldsymbol{z}}$ for each input. Direct encoder optimization requires storing model weights, gradients and intermediate features, leading to prohibitive memory consumption for high-resolution images. We adopt Stochastic Gumbel Annealing (SGA) [80] to refine $\boldsymbol{y}$ and $\boldsymbol{z}$ efficiently. In SGA, outputs of $g_{a}$ and $h_{a}$ serve as initialization. For anchor part $\boldsymbol{y}^{i}_{ac}$, $\boldsymbol{\Lambda}^{i}_{ac}=\boldsymbol{y}^{i}_{ac}-\boldsymbol{\mu}^{i}_{ac}$ is stochastically rounded up or down. The stochastic rounding direction is sampled from a relaxed one-hot categorical distribution:
$\mathit{q}_{\tau}(\boldsymbol{o}^{i}_{ac}|\boldsymbol{\Lambda}^{i}_{ac})\propto\left\{\begin{aligned} \exp\left\{\frac{-\textrm{atanh}(\boldsymbol{\Lambda}^{i}_{ac}-\lfloor\boldsymbol{\Lambda}^{i}_{ac}\rfloor)}{\tau}\right\},\boldsymbol{o}^{i}_{ac}=\{0,1\},\\ \exp\left\{\frac{-\textrm{atanh}(\lceil\boldsymbol{\Lambda}^{i}_{ac}\rceil-\boldsymbol{\Lambda}^{i}_{ac})}{\tau}\right\},\boldsymbol{o}^{i}_{ac}=\{1,0\},\end{aligned}\right$. (14)
where temperature $\tau=\min(0.5,e^{0.001j})$ decreases with refinement step $j$, gradually making the latent representation aware of quantization. Notably, selective compression is not incorporated during SGA. Instead, the selective compression map is predicted by $f_{s}$ based on the overfitted latent representations obtained after SGA. we primarily employ it as a validation tool to demonstrate our model’s potential for instance-level adaptability and performance improvements. Rate control mechanisms [84] could be incorporated in future work for applications requiring strict bit-rate constraints.
Figure 6: PSNR-Bit-rate curve and MS-SSIM-Bit-rate curve. MS-SSIM is converted to dB for better visual clarity. Please zoom in for better view.
   Methods    Venue    BD-Rate (%) w.r.t. VTM 17.0 Intra
   Kodak    Tecnick    CLIC Pro Valid
   PSNR    MS-SSIM    PSNR    PSNR
   Cheng’20 [16]    CVPR’20    5.58    –44.21    7.57    11.71
   ChARM [51]    ICIP’20    3.23    –    –0.88    –
   Qian’21 [55]    ICLR’21    10.05    –39.53    7.52    0.02
   Xie’21 [75]    ACMMM’21    1.55    –43.39    -0.80    3.21
   NLAIC’21 [15]    TIP’21    11.80    –40.27    7.11    –
   Gao’21 [24]    ICCV’21    –8.76    –48.60    -    –6.49
   Guo’22 [26]    TCSVT’22    –4.45    –45.23    –    –
   LBHIC [74]    TCSVT’22    –4.56    –50.54    –    –
   Entroformer [54]    ICLR’22    4.73    –42.64    2.31    –1.04
   SwinT-ChARM [87]    ICLR’22    –1.73    –    –    –
   NeuralSyntax [71]    CVPR’22    8.97    –39.56    –    5.64
   STF [88]    CVPR’22    –2.48    –47.72    –2.75    0.42
   WACNN [88]    CVPR’22    –2.95    –47.71    –5.09    0.04
   ELIC [30]    CVPR’22    –5.95    –44.60    –9.14    –3.45
   Contextformer [38]    ECCV’22    –5.77    –46.12    –9.05    –
   GLLMM [22]    TIP’23    –1.43    –47.41    –5.58    –
   NVTC [21]    CVPR’23    –1.04    –    –    –
   LIC-TCM Large [45]    CVPR’23    –10.14    –48.94    –11.47    –8.04
   QARV [19]    TPAMI’24    0.31    –    –3.03    -
   FTIC [40]    ICLR’24    –12.99    –51.13    –14.88    –9.53
   LLIC-TCM [32]    TMM’24    –10.94    –49.73    –14.99    –10.41
   WeConvene [23]    ECCV’24    –6.71    –48.63    –8.42    –5.74
   CCA [27]    NeurIPS’24    –12.04    –    –14.55    –10.75
   MambaIC [82]    CVPR’25    –13.03    –    –18.41    –15.58
   LALIC [20]    CVPR’25    –13.72    –50.86    –18.19    –14.34
   DCAE [47]    CVPR’25    –15.43    –53.20    –20.51    –15.78
   Baseline Models
   MLIC [34]    ACMMM’23    -8.05    –49.13    –12.73    –8.79
   MLIC+ [34]    ACMMM’23    -11.39    –52.75    –16.38    –12.56
   MLIC++ [33]    ICMLW’23    -13.39    -53.63    -17.59    -13.08
   MLICv2    Ours    -16.54    -54.56    -21.61    -16.05
   MLICv2+    Ours    -20.46    -55.86    -24.35    -19.14
Table 1: BD-Rate $(\%)$ comparison for PSNR (dB) and MS-SSIM. The anchor is VTM-17.0 Intra. “-” means the result is not available.
Figure 7: Reconstructions of our proposed MLICv2, recent learned image compression models [33, 30, 88, 16] and VTM-17.0 Intra [12]. “[MSE]” denotes the model is optimized for MSE. Please zoom in for better view.
(1) In MLICv2+∗, gradient checkpointing [14] is employed during the SGA iterations.
(2) The MACs of MLICv2+ are theoretically estimated based on the number of SGA iterations, following the observation from Epoch AI that backward and forward passes almost always have MAC ratios close to 2:122footnotemark: 2. The MACs of MLICv2+∗ are not reported due to the use of gradient checkpointing, which prevents accurate profiling.
(3) The lack of reported results for some models at certain resolutions is due to out-of-memory (OOM) issues or inaccurate profiling.
Figure 8: GPU memory consumption, Forward MACs, encoding time and decoding time comparisons among our proposed models and recent learned image compression models  [16, 75, 71, 88, 30, 45, 40, 34, 33].
4 Experiments
4.1 Implementation Details
4.1.1 Training Dataset
Our training dataset contains 101,104 images with resolutions exceeding $640\times 640$ pixels. These images are carefully selected from various established image datasets [18, 43, 1, 42, 77]. To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al. [10] by further down-sampling the JPEG images using a randomized factor. Additionally, images with bits per pixel (bpp) lower than 3 are excluded to ensure data quality.
4.1.2 Training Strategy
MLICv2 is implemented using PyTorch 2.2.2 and CompressAI 1.2.6. Following CompressAI, we configure $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE optimization and $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) [73]. Training is conducted on eight Tesla A100-80G GPUs with a batch size of 16, employing the Adam optimizer $\beta_{1}=0.9,\beta_{2}=0.999$. Our training consists of two stages: Stage 1: We train the model with only hyperprior for 2 million (M) steps using a learning rate of $10^{-4}$ and $256\times 256$ patches, providing superior initialization for analysis and synthesis transforms. Stage 2: Loading the pre-trained weights, we train the complete model for 2M steps. The learning rate follows a scheduled decay: starting at $10^{-4}$, then decreasing to $3\times 10^{-5}$, $10^{-5}$, $3\times 10^{-6}$ and $10^{-6}$, at 1.5M, 1.8M, 1.9M, and 1.95M steps, respectively. We employ $256\times 256$ patches for the initial 1.2M steps, then transition to $512\times 512$ patches to leverage the effectiveness of global context modules. For the guided selective compression module optimization, we conduct 20K steps with batch size 16 and learning rate $10^{-3}$.
To validate our model’s potential for further improvement, we employ SGA in MLICv2+, refining latent representations and side information for 3000 steps with learning rate $10^{-3}$. This approach allows us to further enhance the performance under our architectural framework.
4.2 Benchmarks and Metrics
We evaluate rate-distortion performance on three datasets extensively used in learned compression research [9, 64, 10, 50, 51, 16, 75, 31, 30, 88, 87, 38, 19, 71, 26, 74, 15, 45, 34, 33, 32]:
  • Kodak [37]: Contains 24 uncompressed images with $768\times 512$ pixels.
  • Tecnick [4]: Comprises 100 uncompressed images with $1200\times 1200$ pixels.
  • CLIC Pro Valid [65]: The validation set from the 3rd Challenge on Learned Image Compression, consisting of 41 high-resolution images with $2048\times 1440$ pixels.
Performance ranking employs the Bjøntegaard delta rate (BD-Rate) metric [11] for comprehensive evaluation.
4.3 Rate-Distortion Performance
4.3.1 Quantitive Results
For a fair comparison with learned image compression models, VTM-17.0 Intra is evaluated in YUV444 color space under encoder_intra_vtm.cfg. To demonstrate the superiority of our MLICv2 and MLICv2+, we compare with recent learned image compression models [41, 16, 74, 51, 55, 75, 15, 26, 54, 87, 88, 30, 38, 22, 21, 45, 40, 32, 23, 20, 47].
The rate-distortion curves are presented in Fig. 6. When compared with our baseline MLIC++, MLICv2+ performs significantly better, especially at high bit-rates and on high-resolution images [4, 65]. Specifically, our MLICv2 achieves a maximum improvement of $0.6,0.5,0.5$ dB when the bpp is around $0.8,0.55,0.6$ on Kodak, Tecnick, CLIC Pro Val, respectively. Our MLICv2+ achieves larger improvement compared to MLIC++ [33] due to adopted latent refinement [80] for enhanced instance adaptability. The average improvements of latent refinement on MLICv2 on Kodak [37], Tecnick [4], CLIC Pro Val [65] are $0.22,0.15,0.15$ dB, respectively.
BD-Rate reductions over VTM-17.0 Intra are presented in Table 1. Compared with existing methods, our MLICv2 and MLICv2+ achieve state-of-the-art performance, reducing BD-Rates by $16.54\%$ and $20.46\%$ on Kodak. When compared to the baseline model MLIC++ [33], our MLICv2 and MLICv2+ reduce $3.15\%$ and $7.07\%$ more bit-rates over VTM-17.0 Intra on Kodak [37], respectively. Our MLICv2 and MLICv2+ also achieve superior performances on high-resolution datasets Tecnick [4] and CLIC Pro Val [65]. Specifically, our MLICv2 and MLICv2+ reduce BD-Rates by $21.61\%$ and $24.35\%$ on Tecnick [4] and $16.05\%$ and $19.14\%$ on CLIC Pro Val [65].
4.3.2 Qualitative Results
We also compare the qualitative performance to MLIC++, ELIC [30], STF, WACNN [88], Cheng’20 [16] and VTM-17.0 Intra on subjective qualities. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of MLICv2 have sharper textures and retain more details. In terms of visual quality, our MLICv2 has significant improvements on rate-perception performance compared to other models.
4.4 Computational Complexity
The model complexities and sizes are presented in Fig. 8 and Table 5, respectively. To comprehensively evaluate model complexity, 16 images from the LIU4K dataset [44] are cropped to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072\}$ patches to cover the various resolutions of images that can be encountered in practical scenarios. Forward GPU memory, encoding time, decoding time, and forward Multiply-Accumulate operations (MACs) are employed as the complexity metrics. Encoding and decoding times include arithmetic encoding and decoding times. Since the number of skipped elements at different bit-rates varies, we report the results of MLICv2 and MLICv2+ with $\lambda=0.0130$.
Due to the high memory consumption introduced by SGA, we employ gradient checkpointing [14] to overcome GPU memory limitations. This technique alleviates memory pressure by selectively recomputing a subset of intermediate activations during the backward pass. Instead of storing all intermediate activations during the forward pass, only a small set—typically those essential for gradient computation—are cached. The remaining activations are recomputed on-the-fly during backpropagation. This strategy significantly reduces peak memory usage, albeit at the cost of increased computation time.
(a) Kodim01
(b) $j=900$
(c) $j=1101$
(d) Kodim20
(e) $j=1018$
(f) $j=1487$
Figure 9: Visualization of attention maps of hyperprior guided intra slice global context module. The inputs are Kodim01 and Kodim20. The attention map is computed by $\textrm{Softmax}_{2}(\hat{\boldsymbol{H}}_{na,q})[j]\textrm{Softmax}_{1}(\hat{\boldsymbol{H}}_{ac,k})^{\top}$, where $j$ is the query index.
11footnotetext: https://epoch.ai/blog/backward-forward-FLOP-ratio
4.4.1 On Forward GPU Memory
Compared to MLIC+[34], MLICv2 significantly reduces memory usage thanks to its linear-complexity entropy model. Relative to MLIC++[33], it incurs slightly higher memory for high-resolution images due to the additional modules introduced in this work, but still requires less memory than Xie’21 [75], Qian’21 [55], Entroformer [55], and LIC-TCM [45]. This modest increase is justified by the corresponding performance improvements. Latent refinement, while increasing memory during backpropagation, is optional and primarily benefits scenarios where encoding complexity is less critical. Since images are typically encoded once but decoded multiple times, the added encoding cost is acceptable given the notable gains in rate-distortion performance.
Figure 10: Upper-bound reconstruction quality (PSNR) on the Kodak dataset without entropy constraints.
Transform Architecture Cheng’20 [16] SwinT [87] ELIC [30] MLIC++ LIC-TCM [45] MLICv2
BD-Rate(%) 0.00 -4.61 -5.19 -0.41 -8.91 -9.44
Table 2: Comparisons among different transforms of proposed MLICv2, LIC-TCM [45], MLIC++, ELIC [30], and Cheng’20 [16]. All method are quiped with autoregressive context module [52].
Input Image Selected Channel Features of Intra-Slice Local Context and Mean Attention Weights
Kodak: Kodim20 Mean Attention Weight: 0.1555 Mean Attention Weight: 0.1211 Mean Attention Weight: 0.0553 Mean Attention Weight: 0.0127
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kodak: Kodim21 Mean Attention Weight: 0.1471 Mean Attention Weight: 0.0780 Mean Attention Weight: 0.0401 Mean Attention Weight: 0.0184
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Image Selected Channel Features of Inter-Slice Local Context and Mean Attention Weights
Kodak: Kodim20 Mean Attention Weight: 0.0933 Mean Attention Weight: 0.0822 Mean Attention Weight: 0.0503 Mean Attention Weight: 0.0361
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kodak: Kodim21 Mean Attention Weight: 0.0909 Mean Attention Weight: 0.0720 Mean Attention Weight: 0.0452 Mean Attention Weight: 0.0327
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Image Selected Channel Features of Intra-Slice Global Context and Mean Attention Weights
Kodak: Kodim20 Mean Attention Weight: 0.0728 Mean Attention Weight: 0.0704 Mean Attention Weight: 0.0619 Mean Attention Weight: 0.0393
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kodak: Kodim21 Mean Attention Weight: 0.0648 Mean Attention Weight: 0.0623 Mean Attention Weight: 0.0607 Mean Attention Weight: 0.0234
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Image Selected Channel Features of Inter-Slice Global Context and Mean Attention Weights
Kodak: Kodim20 Mean Attention Weight: 0.1424 Mean Attention Weight: 0.1378 Mean Attention Weight: 0.0783 Mean Attention Weight: 0.0134
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kodak: Kodim21 Mean Attention Weight: 0.2302 Mean Attention Weight: 0.1499 Mean Attention Weight: 0.1391 Mean Attention Weight: 0.0107
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Visualization of selected channel features and their corresponding Mean Attention Weights. All channels are sampled from the context features (before context reweighting) of the last slice. The visualized feature maps are normalized to the range [0, 1].
Context Module MLIC++ MLICv2 Hyperprior Guided Global Context 2D RoPE Context Reweighting Selective Compression
GMACs $508.49$ $506.76$ $1.62$ $2.4\times 10^{-5}$ $3.46$ $1.73$
Table 3: MACs contribution of the introduced modules to the entropy model on the Kodak dataset
4.4.2 On Forward MACs
Compared to MLIC+, LIC-TCM [45], our MLICv2 has lower forward MACs. Compared to MLIC++, our MLICv2 exhibits slightly higher forward MACs. This increase can be attributed to two primary factors: the proposed channel reweighting module and selective compression module. However, considering the significant performance gains achieved through these enhancements, the marginal increase in MACs is deemed acceptable. The trade-off between computational complexity and improved compression efficiency is well-balanced in our proposed model.
4.4.3 On Encoding Time and Decoding Time
Compared to MLIC and MLIC+ [34], MLICv2 achieves higher encoding and decoding efficiency due to its linear-complexity components. Relative to MLIC++[33], it further reduces encoding and decoding time by efficiently skipping many zero elements, which is particularly beneficial for high-resolution images where zeros are more frequent. Latent refinement, while involving multiple backpropagation steps and increasing encoding time, is optional and suited for scenarios where encoding complexity is less critical. Since images are typically encoded once but decoded multiple times, the extra encoding cost is justified by the performance gains. Notably, MLICv2+ maintains decoding complexity nearly identical to MLICv2, with both substantially faster than MLIC++, recent WeConvene [23] and DCAE [47].
4.4.4 On Model Size
The model size comparisons are presented in Table 5. Although MLICv2 incorporates several additional components, its parameter count (84.3M) remains comparable to that of MLIC++ (83.5M). The increase in parameters is minimal, primarily because we deliberately reduce the number of channels for intermediate features in the context modules of MLICv2 to control overall model size.
Context Modules MLICv2+ Cumulative Benefits Individual Contributions
Full Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8 Case 9 Case 10 Case 11 Case 12 Case 13 Case 14 Case 15 Case 16
Latent Refinement
Selective Compression
2D RoPE
Context Reweighting
Hyperprior Guided Global
Simple Token Mixing Transform
Gate
BD-Rate (%) 0.00 0.17 4.89 5.21 5.77 7.79 8.29 9.58 9.87 4.64 9.52 9.12 7.64 8.77 8.51 9.58 9.87
Table 4: Complete ablation study on Kodak dataset. BD-Rate (%) is employed to evaluate component contributions. MLICv2+ is the anchor. Cases 1-8 show cumulative benefits, Cases 9-16 show individual contributions.
4.5 Ablation Studies
4.5.1 Settings
Each ablation cases is optimized for MSE using our established training strategy, with MLICv2+ serving as the anchor. We evaluate both cumulative benefits and individual contributions of each proposed component. The results are presented in Table 4.
Figure 12: Distribution of scale values in MLICv2 on Kodak [37].
4.5.2 Analysis of Enhanced Transform
To demonstrate the superior capacity of our proposed transform, we compare the upper bounds of reconstruction quality achieved by the transforms in MLIC++[33], LIC-TCM[45], and our MLICv2. The results are shown in Fig. 10. Our MetaFormer-based design achieves better reconstruction quality than MLIC++ and performs comparably to LIC-TCM, highlighting its stronger representational power despite relying primarily on convolutional operations.
To demonstrate the effectiveness of the proposed simple token mixing block, we compare the transform module with the transform modules of Cheng’20 [16], SwinT-Charm [87], ELIC [30], and LIC-TCM [45] on rate-distortion performance. The entropy model is aligned with spatial autoregressive entropy model [50]. The results are presented in Table. 2. Our proposed transform achieves better performance than LIC-TCM while with about half forward MACs. When equiped with proposed transform, the model performance gets further enhanced. Since our baseline performance is very strong, the boost from the transform is slightly small relative to Table. 2.
4.5.3 Analysis of Enhanced Entropy Modeling
As shown in Table 4, disabling the selective compression module results in a 0.17% performance drop, which is acceptable, as the skipped elements—primarily zeros—typically have very low entropy despite accounting for the majority of all latent elements. By omitting the encoding of these zero elements, we can significantly accelerate both arithmetic encoding and decoding. To demonstrate the effectiveness of proposed guided selective compression on acceleration, the skip ratio, encoding acceleration and decoding acceleration on Kodak [37] and CLIC Pro Val [65] are presented in Table 6. As image quality improves, the proportion of zero elements that can be skipped decreases, which in turn reduces the acceleration gains for both encoding and decoding. However, high-resolution images usually contain more redundancy, allowing a larger number of elements to be skipped. This leads to greater potential for acceleration in such cases. In addition, we found that setting $\xi$ to 0.2, 0.3, or 0.4 has little effect on overall performance. This is mainly due to two reasons. First, as illustrated in Figure 12, most scale values fall within the range [0.1, 0.2], with only a small fraction between 0.2 and 0.4. Thus, using $\xi$ within this range is already effective in skipping most zero elements with low prediction error. The selective compression module only needs to correct a few wrongly skipped important elements. Second, we apply a weighted binary classification loss to train the selective compression module, using the ground-truth encoding mask as supervision.
To demonstrate the effectiveness of hyperprior-guided global context for the first slice, the attention maps are illustrated in Fig. 9, which are generally accurate, thereby enhancing the performance. For slices $\hat{\boldsymbol{y}}^{\geq 1}$, the previous one slice is still employed to predict the global correlations due to the higher bit-rate of each slice compared to the side information.
To demonstrate the relationship between context reweighting map and image content, we visualize the Mean Attention Weight in Fig. 11.
Model
Xie’21 [75]
ELIC [30]
WACNN [88]
STF [88]
LIC-TCM [45]
FTIC [40]
MLIC+ [34]
MLIC++ [33]
WeConvene [23]
LALIC [20]
DCAE [47]
MLICv2
Parameters (M) 50.0 41.9 75.2 100.0 75.9 69.8 83.4 83.5 105.5 63.2 119.2 84.3
Table 5: Model Size Comparison.
Kodak CLIC Pro Val
$\lambda$ 0.0018 0.0035 0.0067 0.0130 0.0250 0.0483 0.0018 0.0035 0.0067 0.0130 0.0250 0.0483
Skip Ratio (%) 88.32 82.73 77.09 69.76 66.44 57.26 92.95 88.20 83.72 79.72 77.28 68.81
Enc Acceleration (%) 9.98 9.48 9.24 8.81 7.31 7.22 18.48 17.12 16.88 16.62 16.15 14.21
Dec Acceleration (%) 13.61 13.21 12.15 11.97 10.18 9.62 27.52 25.85 25.37 24.21 23.36 20.99
Table 6: Skip Ratio, Encoding Acceleration and Decoding Acceleration at different qualities on Kodak and CLIC Pro Val.
Figure 13: Changes in loss, PSNR and bpp during the refinement of latent representation and side information of Kodim01 from Kodak dataset [37]. “Initial” denotes the original loss, psnr, bpp when using the latent representation and side information outputed by MLICv2.
Definition 4.1 (Mean Attention Weight).
In the channel reweighting module, the attention map is computed as $\boldsymbol{M}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})\in\mathbb{R}^{C\times C}$, where $C$ is the number of channels. The attention map is row-wise normalized via the Softmax function: $\boldsymbol{M}_{i,j}=\frac{\exp(\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j})}{\sum_{j^{\ast}=1}^{C}\exp(\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j^{\ast}})},$ where $\boldsymbol{q}_{i}$ and $\boldsymbol{k}_{j}$ are the $i$-th and $j$-th channel vectors of $\boldsymbol{Q}$ and $\boldsymbol{K}$, respectively. This normalization ensures that each row of $\boldsymbol{M}$ sums to 1, meaning that $\boldsymbol{M}_{i,j}$ reflects the relative influence of input channel $j$ on output channel $i$ when computing the weighted sum: $\boldsymbol{o}_{i}=\sum_{j=1}^{C}\boldsymbol{M}_{i,j}\boldsymbol{v}_{j},$ where $\boldsymbol{v}_{j}$ is $j$-th channel vector $\boldsymbol{V}$. The Mean Attention Weight for each input channel $j$ is then defined as the average over the $j$-th column of $\boldsymbol{M}$: $\tilde{\boldsymbol{M}}_{j}=\frac{1}{C}\sum_{i=1}^{C}\boldsymbol{M}_{i,j},$ which reflects the overall influence of input channel $j$ across all output channels.
As observed, channels dominated by noise tend to receive lower Mean Attention Weight, while those containing richer structural information or less noise are assigned higher weight. This demonstrates that our context reweighting module effectively emphasizes more informative contexts, supporting its contribution to the performance gains.
To validate the effectiveness of the proposed 2D RoPE, the 2D RoPE is replaced by original additive relative position embedding [46]. The performance degradation caused by the replacement demonstrates the superiority of the proposed 2D RoPE. Inspired by some 2D positional encoding designs [13, 57, 72], which split the feature tensor along the channel dimension and apply positional encoding along the horizontal axis to one half and the vertical axis to the other, we also experimented with a similar strategy. Specifically, we divided the input tensor along the channel dimension into two parts and applied RoPE-based rotations along the horizontal axis and vertical axis, respectively. We found that this alternative design does not lead to a performance difference. We also conduct experiments on the baseline model using 2D RoPE without learnable $\theta_{x}$ and $\theta_{y}$. This variant results in approximately a 0.4% performance drop.
The validate the effectiveness of the gate block, the gate block is replaced by Multilayer Perceptron (MLP), which leads to performance loss.
Overall, the newly introduced modules in the entropy model yield substantial performance gains with minimal additional computational cost, as demonstrated in our ablation studies and Table 3. As reported in Table 3, the MACs of 2D RoPE are almost negligible. The context reweighting and selective compression modules contribute only $0.68\%$ and $0.34\%$ to the total MACs on the Kodak dataset, respectively. This efficiency is largely due to the fact that the proposed modules for entropy modeling operate in the latent space, where the spatial resolution is only $\bm{\frac{1}{16}\times\frac{1}{16}}$ of the input image. Such a drastic reduction in resolution greatly reduces the overall computational burden.
4.5.4 Analysis of Iterative Latent Refinement
In case 1, the latent refinement is removed, which leads to around $4.89\%$ bit-rate increase compared to MLICv2+. The increased bit-rate demonstrates the effectiveness of the latent refinement for exploring performance potential. The changes of loss, PSNR and bpp are illustrated in Fig. 13. During the refinement, the loss and bpp generally first rise and then fall to lower than the initial loss and bpp, while the PSNR generally keeps rising.
5 Conclusion
In this paper, we present MLICv2 and MLICv2+, enhanced versions of the previous MLIC series that address observed limitations in transform design and entropy modeling. To overcome these limitations, we introduce efficient simple token mixing blocks that substantially improve representational capacity while maintaining computational efficiency. Our MetaFormer-based design achieves comparable performance to complex transformer alternatives with reduced overhead. For entropy modeling, we propose hyperprior-guided global context prediction, context reweighting mechanisms, and novel 2D Rotary Positional Embedding, collectively providing richer contextual information. The guided selective compression further accelerates encoding/decoding through intelligent zero-element skipping. Additionally, by employing SGA, we demonstrate the potential for performance improvements within our architectural framework. This validation analysis reveals the potential for further architectural improvements and provides valuable insights for future development directions. These innovations collectively establish new state-of-the-art performance benchmarks, with BD-Rate reductions of up to 20% compared to VTM-17.0 Intra while maintaining competitive computational complexity. Furthermore, to continue exploring the performance upper bounds via encoding-time scaling, we plan to investigate the transmission of update parameters [67, 59, 49] to the decoder side.
Limitations: Although our MLICv2 and MLICv2+ achieve state-of-the-art performance, several limitations warrant future attention. First, the performance potential exploration through SGA may lead to bitrate variations [84], requiring practical online encoder optimization strategies for precise bitrate control. Second, MLICv2/MLICv2+ may encounter cross-platform decoding issues [7, 29, 61, 79]. We plan to investigate model quantization techniques [28, 29, 61] and chain coding-based latent compression [79] to address deployment challenges.

### v3

@@
 Notably, the MLIC series—LICs equipped with multi-reference entropy models—have substantially surpassed conventional image codecs such as Versatile Video Coding (VVC) Intra.
-However, existing MLIC variants suffer from several limitations: performance degradation at high bitrates due to insufficient transform capacity, suboptimal entropy modeling that fails to capture global correlations in initial slices, and lack of adaptive channel importance modeling.
+However, existing MLIC variants suffer from several limitations: performance degradation at high bit-rates due to insufficient transform capacity, suboptimal entropy modeling that fails to capture global correlations in initial slices, and lack of adaptive channel importance modeling.
 In this paper, we propose MLICv2 and MLICv2+, enhanced successors that systematically address these limitations through improved transform design, advanced entropy modeling, and exploration of the potential of instance-specific optimization.
-For transform enhancement, we introduce a lightweight token mixing block inspired by the MetaFormer architecture, which effectively mitigates high-bitrate performance degradation while maintaining computational efficiency.
+For transform enhancement, we introduce a lightweight token mixing block inspired by the MetaFormer architecture, which effectively mitigates high-bit-rate performance degradation while maintaining computational efficiency.
 For entropy modeling improvements, we propose hyperprior-guided global correlation prediction to extract global context even in the initial slice of latent representation, complemented by a channel reweighting module that dynamically emphasizes informative channels.
@@
 1 Introduction
-Image compression constitutes a fundamental challenge in multimedia communication, with the objective of achieving compact representation while preserving reconstruction quality.
-Traditional compression methods, exemplified by VVC Intra [12], rely on block-based hybrid coding frameworks refined over decades.
-However, as conventional approaches near performance saturation, learned image compression (LIC) has emerged as a transformative paradigm, demonstrating substantial potential to surpass hand-crafted codecs through end-to-end optimization.
-Contemporary LIC methods [9, 64, 10, 50, 51, 16, 75, 31, 30, 88, 35, 86, 68, 87, 38, 19, 71, 26, 74, 78, 53, 15, 45, 34, 33, 32] predominantly adopt autoencoder-based frameworks with rate-distortion optimization, comprising analysis transforms, quantization, synthesis transforms, and entropy models.
-The analysis transform maps input images to latent representations, which undergo quantization before reconstruction via synthesis transforms.
-Entropy models estimate quantized latent distributions to facilitate efficient entropy coding.
-Advanced techniques of transform architectures and entropy models have enabled several LICs [30, 34, 33, 32, 45, 40, 47, 20] to achieve performance comparable to or exceeding advanced traditional codecs like VVC Intra [12].
-Figure 1: BD-Rate and decoding time during inference on CLIC Professional Valid [65] with 2K resolution. The bubble size indicates the number of model parameters.
-Entropy modeling plays a pivotal role in LIC performance, aiming to minimize conditional entropy through informative prior capture.
-While hyperprior models [10] provide side information, context models [52, 31] capture spatial dependencies between previously decoded and current elements.
-Building on these foundations, Jiang et al.
-introduced the MLIC series [34, 33], which leverages multi-reference priors through latent representation partitioning into channel-wise slices [51], capturing both inter-slice and intra-slice dependencies.
-MLIC employs checkerboard partitioning [31] within slices for two-pass decoding, while attention mechanisms model global dependencies from previous slices.
-MLIC++ [33] addressed computational complexity by replacing vanilla attention [70] with linear attention [60] and modeling global correlations across previous slices.
-Despite the strong overall performance of the MLIC series, several critical limitations persist and motivate our work.
-First, regarding transform design, MLIC models exhibit performance degradation at high bitrates, primarily attributed to insufficient transform capacity for preserving latent representation information.
-Second, concerning entropy modeling, global correlation estimation relies exclusively on previous slices, precluding global correlation capture in initial slices.
-Moreover, existing approaches treat all channels equally without considering varying informational importance, potentially limiting entropy modeling accuracy.
-To systematically address these limitations, we propose MLICv2, featuring enhanced transform and entropy modeling modules, and MLICv2+, which additionally incorporates iterative refinement techniques [80] to demonstrate framework potential.
-For transform improvement, we design an efficient token mixing block replacing residual blocks in MLIC++.
-Inspired by MetaFormer architecture [81], our block employs depth-wise convolution for spatial token mixing and lightweight gated interaction for point-wise operations, achieving superior high-bitrate performance with reduced complexity.
-For entropy modeling enhancement, we propose MEMv2, an advanced multi-reference entropy model capturing diverse dependencies.
-We introduce hyperprior-guided global correlation prediction for initial slices, leveraging spatial distribution similarity between hyperprior and latent representations.
-Our two-stage context modeling strategy performs spatial aggregation followed by adaptive channel-wise reweighting to capture feature importance variations.
-Additional improvements include 2D Rotary Positional Embedding (RoPE) and guided selective compression for enhanced modeling and coding efficiency.
-Additionally, we employ Stochastic Gumbel Annealing (SGA) [80] to iteratively refine latent representations and side information for each input.
-This approach demonstrates our framework’s potential for input-specific rate-distortion optimization while preserving decoder complexity, suggesting opportunities for further architectural enhancements.
-Our comprehensive approach yields significant improvements, with MLICv2 and MLICv2+ substantially outperforming MLIC++ as demonstrated in Figure 1.
-MLICv2 achieves BD-rate reductions of 16.54%, 21.61%, and 16.05% on Kodak, Tecnick, and CLIC Pro Val datasets respectively, while MLICv2+ further improves these results to 20.46%, 24.35%, and 19.14% over VTM-17.0 Intra.
+Image compression is a fundamental problem in multimedia communication, aiming to achieve compact representation while maintaining high reconstruction quality.
+Traditional codecs such as VVC Intra bross2021vvc rely on block-based hybrid coding frameworks refined over decades.
+However, as these hand-crafted methods approach performance saturation, learned image compression (LIC) has emerged as a new paradigm that leverages end-to-end optimization to outperform conventional codecs.
+Recent LIC approaches balle2016end, theis2017lossy, balle2018variational, minnen2018joint, minnen2020channel, cheng2020learned, xie2021enhanced, he2021checkerboard, he2022elic, zou2022the, jin2025neural, zhao2023universal, tu2024reconstruction, zhu2022transformerbased, koyuncu2022contextformer, duan2023qarv, wang2022neural, guo2021causal, wu22block, yang2024perceptual, mishra2021multi, mishra2020wavelet, mishra2022deep2, mishra2022deep, park2023scalable, chen2021nic, liu2023learned, jiang2022mlic, jiang2023mlicpp, jiang2024llic typically follow an autoencoder-based framework optimized for the rate-distortion trade-off.
+An analysis transform encodes the image into latent representations, which are quantized and decoded by a synthesis transform.
+Entropy models then estimate the probability distribution of quantized latents to enable efficient coding.
+With advanced transform designs and entropy models, several LIC systems he2022elic, jiang2022mlic, jiang2023mlicpp, jiang2024llic, liu2023learned, li2024frequencyaware, lu2025learned, feng2025linear have achieved or even surpassed the performance of state-of-the-art traditional codecs such as VVC Intra bross2021vvc.
+Entropy modeling is crucial to LIC performance as it minimizes conditional entropy by capturing informative priors.
+Hyperprior models balle2018variational introduce side information, while context models minnen2018image, he2021checkerboard capture spatial dependencies among neighboring elements.
+Building upon these foundations, Jiang et al.
+proposed the MLIC series jiang2022mlic, jiang2023mlicpp,
+Figure 1: BD-Rate and decoding time during inference on CLIC Professional Valid CLIC2020 with 2K resolution. The bubble size indicates the number of model parameters.
+which partitions latents into channel-wise slices minnen2020channel to model both inter- and intra-slice local111“Local” denotes neighboring regions, specifically within the receptive field of a convolution or a windowed attention.
+and global222“Global” refers to the entire feature map.
+dependencies.
+Within each slice, checkerboard partitioning he2021checkerboard enables two-pass decoding, and attention modules model global correlations from previous slices.
+To reduce computational cost, MLIC++ jiang2023mlicpp replaces standard attention vas2017attention with linear attention shen2021efficient333“Linear” denotes complexity linear with spatial resolution., which scales more efficiently while maintaining global correlation modeling.
+Despite their strong performance, MLIC models still face several limitations that motivate this work.
+First, in terms of transform design, performance degrades at high bit-rates due to limited transform capacity, which restricts the preservation of latent information.
+Second, for entropy modeling, global correlation estimation relies solely on previously decoded slices, preventing the capture of global dependencies in early slices.
+Furthermore, all channels are treated equally without considering their varying informational importance, potentially reducing modeling accuracy.
+To systematically address these limitations, and as illustrated in Fig 2, we propose MLICv2, featuring enhanced transform and entropy modeling modules, and MLICv2+, which additionally incorporates iterative refinement techniques yang2020improving to demonstrate framework potential.
+For transform improvement, we design an efficient token mixing block to replace the residual blocks in MLIC++.
+Inspired by the MetaFormer architecture yu2022metaformer, this block employs depth-wise convolution for spatial mixing and lightweight gated interaction for point-wise operations, achieving better high-bit-rate performance with lower complexity.
+For entropy modeling, we propose MEMv2, a multi-reference entropy model that captures richer dependencies.
+We introduce hyperprior-guided global correlation prediction for initial slices by exploiting the spatial similarity between hyperprior and latent features.
+A two-stage context modeling strategy is also proposed: it first performs spatial aggregation and then applies adaptive channel-wise reweighting to capture feature importance variations.
+Further enhancements include 2D Rotary Positional Embedding and guided selective compression for improved modeling and coding efficiency.
+Finally, we employ Stochastic Gumbel Annealing (SGA) yang2020improving to iteratively refine latent representations and side information for each input, enabling input-specific rate-distortion optimization without increasing decoder complexity.
+This demonstrates the scalability and adaptability of our framework for future extensions.
+Our comprehensive approach leads to significant performance gains, as shown in Fig.
+1.
+MLICv2 achieves BD-rate reductions of 16.54%, 21.61%, and 16.05% on the Kodak, Tecnick, and CLIC Professional Validation datasets, respectively, while MLICv2+ further improves these results to 20.46%, 24.35%, and 19.14% over VTM-17.0 Intra.
 Our contributions are summarized as follows:
-  • We introduce MetaFormer-based architecture for LIC through our simple token mixing block, achieving superior high-bitrate performance with reduced computational complexity compared to previous residual blocks.
-  • We systematically improve entropy modeling through multiple improvements: (i) hyperprior-guided global correlation prediction addressing initial slice limitations, (ii) the first two-stage context modeling strategy with spatial aggregation and adaptive channel reweighting, and (iii) advanced positional embedding and selective compression mechanisms.
-  • We demonstrate the effectiveness of our framework by applying iterative refinement techniques (SGA), showing consistent performance gains without decoder complexity increase and suggesting opportunities for further architectural enhancements.
-  • Extensive experiments validate our approach, with MLICv2 and MLICv2+ achieving BD-rate reductions of 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% over VTM-17.0 Intra on Kodak, Tecnick, and CLIC Pro Val datasets respectively.
+Figure 2: Model evolution of the MLIC series.
+  • We introduce a MetaFormer-based transform architecture for LIC through an efficient token mixing block, achieving better high-bit-rate performance and lower complexity than previous residual designs.
+  • We enhance entropy modeling with three key innovations: (i) hyperprior-guided global correlation prediction for initial slices, (ii) the first two-stage context modeling strategy with spatial aggregation and adaptive channel reweighting, and (iii) advanced positional embedding and selective compression mechanisms.
+  • We apply iterative refinement via SGA to demonstrate consistent rate-distortion improvements without increasing decoder complexity.
+  • Extensive experiments validate the effectiveness of MLICv2 and MLICv2+, achieving BD-rate reductions of 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% on the Kodak, Tecnick, and CLIC Professional Validation datasets, respectively.
 2 Related Works
 2.1 Learned Image Compression
-Autoencoder-based learned image compression [9, 64, 50, 16, 30, 34, 33, 51, 32] follows the transform coding paradigm, consisting of an analysis transform $g_{a}$ with parameters $\phi$, a synthesis transform $g_{s}$ with parameters $\psi$, a quantization function $\lceil\cdot\rfloor$, and an entropy model $p$ with parameters $\rho$.
-Given an input image $\boldsymbol{x}\sim p_{\boldsymbol{x}}$, the process is formulated as:
-$\boldsymbol{y}=g_{a}(\boldsymbol{x};\phi),\hat{\boldsymbol{y}}=\lceil\boldsymbol{y}\rfloor,\hat{\boldsymbol{x}}=g_{s}(\hat{\boldsymbol{y}};\psi).$ (1)
-The rate-distortion optimization is achieved through:
-$\mathcal{L}=\mathbb{E}_{\boldsymbol{x}\sim p_{\boldsymbol{x}}}\left[\underbrace{-\log_{2}\int p(\hat{\boldsymbol{y}};\rho)d\hat{\boldsymbol{y}}}_{\textrm{bit-rate}}+\lambda\times\underbrace{\mathcal{D}(\hat{\boldsymbol{x}},\boldsymbol{x})}_{\textrm{distortion}}\right],$ (2)
-where $\lambda$ balances the rate-distortion trade-off.
-During training, quantization is approximated by adding uniform noise [9] or straight-through estimation [64].
-Ballé et al.
-[10] introduce hyperprior networks using hyper analysis $h_{a}$ and hyper synthesis $h_{s}$ to improve entropy estimation through quantized side information $\hat{\boldsymbol{z}}$ extracted from latent representations.
-Transform architectures have evolved significantly, incorporating Generalized Divisive Normalization (GDN) [8], residual non-local attention [85, 15], deep residual networks [16], and transformer-based designs [87, 88, 48].
-Mixed CNN-transformer [45] and Mamba [82, 25, 56] architectures have also been explored for capturing both local and non-local interactions.
-Entropy estimation has progressed from simple univariate Gaussian models [10] to sophisticated approaches including mean-scale Gaussian models [50], asymmetric Gaussian distributions [17], Gaussian mixture models [16], and generalized Gaussian models [83], significantly improving compression efficiency.
+Autoencoder-based learned image compression balle2016end, theis2017lossy, minnen2018joint, cheng2020learned, he2022elic, jiang2022mlic, jiang2023mlicpp, minnen2020channel, jiang2024llic operates within the transform coding framework, where an input image $\boldsymbol{x}\sim p_{\boldsymbol{x}}$ is first mapped to a latent representation $\boldsymbol{y}=g_{a}(\boldsymbol{x};\phi)$ by an analysis transform $g_{a}$ with parameters $\phi$, then quantized as $\hat{\boldsymbol{y}}=\lceil\boldsymbol{y}\rfloor$ and finally reconstructed as $\hat{\boldsymbol{x}}=g_{s}(\hat{\boldsymbol{y}};\psi)$ using a synthesis transform $g_{s}$ with parameters $\psi$.
+A parametric entropy model $p(\cdot;\rho)$ with parameters $\rho$ is employed to estimate the latent distribution for efficient compression.
+During training, quantization is approximated by adding uniform noise $\boldsymbol{u}\sim\mathcal{U}(-0.5,0.5)$ or by using the straight-through estimator (STE) theis2017lossy.
+Following Ballé et al.
+balle2018variational, the training objective is formulated as:
+$\mathcal{L}=\mathcal{R}(\hat{\boldsymbol{y}})+\lambda\,\mathcal{D}(\hat{\boldsymbol{x}},\boldsymbol{x})$}, (1)
+[…]
+Transform architectures have evolved significantly, incorporating Generalized Divisive Normalization (GDN) balle2015gdn, wavelets fu2024weconvene, mishra2020wavelet, attentions zhangresidual, chen2021nic, mishra2022deep2, multi-scale networks mishra2022deep, deep residual networks cheng2020learned, and transformer-based designs zhu2022transformerbased, zou2022the, lu2021tic.
+Mixed CNN-transformer liu2023learned and Mamba zeng2025mambaic, gumamba, qin2024mambavc architectures have also been explored for capturing both local and non-local interactions.
+Entropy estimation has progressed from simple univariate Gaussian models balle2018variational to sophisticated approaches including mean-scale Gaussian models minnen2018joint, asymmetric Gaussian distributions cui2021asym, Gaussian mixture models cheng2020learned, and generalized Gaussian models zhang2025generalized, significantly improving compression efficiency.
 2.2 Conditional Entropy Models
@@
 […]
-Local context has evolved from serial approaches like PixelCNN [69, 50] to efficient parallel methods including zigzag scanning [41] and checkerboard partitioning [30].
-These intra-slice local contexts are complemented by inter-slice approaches and channel-wise autoregressive coding [51] for faster decoding.
-Recent works investigate long-range correlations among distant elements [55, 26, 36, 34, 33].
-Qian et al.
-[55, 54] employ neighbor elements and bi-directional attention for global correlation estimation.
-Guo et al.
-[26] utilize L-2 distance metrics, while Kim et al.
-[36] transmit global information to decoders.
+Ballé et al.
+balle2018variational introduce hyperprior networks using hyper analysis $h_{a}$ and hyper synthesis $h_{s}$ to improve entropy estimation through quantized side information $\hat{\boldsymbol{z}}$ extracted from latent representations.
+Local context models have evolved from serial methods, such as PixelCNN van2016conditional, minnen2018joint, to more efficient parallel approaches, including zigzag scanning li2020efficient and checkerboard partitioning he2022elic, which focus on intra-slice dependencies by treating the latent representation as a slice.
+Minnen et al.
+minnen2020channel introduced channel-wise autoregressive coding to efficiently model dependencies along the channel dimension and achieve faster decoding.
+Recent studies further extend context modeling to capture long-range dependencies qian2020learning, guo2021causal, kim2022joint, jiang2022mlic, jiang2023mlicpp.
+For instance, Qian et al.
+qian2020learning, qian2022entroformer introduce bi-directional attention to model global correlations, while Guo et al.
+guo2021causal and Kim et al.
+kim2022joint transmit global priors directly to the decoder for enhanced reconstruction.
 Recently, Jiang et al.
-[34] propose MLIC and MLIC+, dividing global contexts into inter-slice and intra-slice components.
+jiang2022mlic propose MLIC and MLIC+, dividing global contexts into inter-slice and intra-slice components.
 The latent representation is partitioned into multiple slices for autoregressive conditional coding, employing checkerboard attention for intra-slice contexts and convolutional layers for inter-slice contexts.
-However, the quadratic complexity of vanilla attention [70] limits scalability for high-resolution images, which is addressed by MLIC++ [33] through linear attention mechanisms [60].
+However, the quadratic complexity of vanilla attention vas2017attention limits scalability for high-resolution images, which is addressed by MLIC++ jiang2023mlicpp through linear attention mechanisms shen2021efficient.
 Figure 3: The overall architecture of MLICv2/MLICv2+.
@@
 $f_{s}$ is the selective compression predictive module.
-$g_{lrp}$ is the latent residual prediction module [51].
+$g_{lrp}$ is the latent residual prediction module minnen2020channel.
 AE and AD are arithmetic encoding and decoding.
@@
 3.1 Background: MLIC++
-Our proposed MLICv2 and MLICv2+ build upon the foundation established by MLIC++ [33].
-MLIC++ employs a simplified version of Cheng’20’s transform [16], where attention modules are removed to reduce computational complexity.
-For entropy modeling, MLIC++ introduces a linear-complexity multi-reference entropy model (MEM++), which effectively captures hyperprior, intra-slice, and inter-slice dependencies.
+Our proposed MLICv2 and MLICv2+ are built upon the foundation of MLIC++ jiang2023mlicpp.
+MLIC++ adopts a simplified variant of Cheng’20’s transform cheng2020learned, where attention modules are removed to reduce computational complexity.
+For entropy modeling, MLIC++ introduces a linear-complexity multi-reference entropy model (MEM++) that effectively captures hyperprior, intra-slice, and inter-slice dependencies.
 The hyperprior information $\hat{\boldsymbol{H}}$ is derived from side information $\hat{\boldsymbol{z}}$.
 Following the approach of Minnen et al.
-[51], the latent representation $\hat{\boldsymbol{y}}$ is partitioned into $L+1$ slices: $\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\ldots,\hat{\boldsymbol{y}}^{L}$}.
-Each slice $\hat{\boldsymbol{y}}^{i}$ is further divided into anchor $\hat{\boldsymbol{y}}^{i}_{ac}$ and non-anchor $\hat{\boldsymbol{y}}^{i}_{na}$ components using checkerboard partitioning [31].
-The inter-slice local context module $g_{\ell}^{inter}$ captures inter-slice local context $\hat{\boldsymbol{C}}^{<i}_{\ell}$ from preceding slices $\hat{\boldsymbol{y}}^{<i}$.
-To model inter-slice global context $\hat{\boldsymbol{C}}^{<i}_{g}$, MLIC++ employs linear attention [60] to achieve an optimal balance between performance and computational complexity.
+minnen2020channel, the latent representation $\hat{\boldsymbol{y}}$ is partitioned into $L+1$ slices: $\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\ldots,\hat{\boldsymbol{y}}^{L}$}.
+Each slice $\hat{\boldsymbol{y}}^{i}$ is further divided into anchor $\hat{\boldsymbol{y}}^{i}_{ac}$ and non-anchor $\hat{\boldsymbol{y}}^{i}_{na}$ components using checkerboard partitioning he2021checkerboard.
+The inter-slice local context module $g_{\ell}^{inter}$ extracts inter-slice local context $\hat{\boldsymbol{C}}^{<i}_{\ell}$ from preceding slices $\hat{\boldsymbol{y}}^{<i}$.
+To model inter-slice global context $\hat{\boldsymbol{C}}^{<i}_{g}$, MLIC++ employs linear attention shen2021efficient to achieve an optimal balance between performance and computational complexity.
 The intra-slice local context $\hat{\boldsymbol{C}}^{i}_{\ell}$ is extracted from the anchor partition $\hat{\boldsymbol{y}}^{i}_{ac}$ using overlapped window-based checkerboard attention $g_{\ell}^{intra}$.
@@
 The estimated means $\boldsymbol{\mu}^{i}_{ac}$, $\boldsymbol{\mu}^{i}_{na}$ and scales $\boldsymbol{\sigma}^{i}_{ac}$, $\boldsymbol{\sigma}^{i}_{na}$ are extracted from the contextual information using the entropy parameter module $g_{ep}$.
-MLIC++ incorporates mixed quantization [51] and the latent residual prediction module $g_{lrp}$ [51].
-In mixed quantization, zero-centered straight-through estimation (STE) [64] is applied for distortion computation, while additive uniform noise (AUN) [9] is used for entropy estimation during training.
+MLIC++ incorporates mixed quantization minnen2020channel and the latent residual prediction module $g_{lrp}$ minnen2020channel.
+In mixed quantization, zero-centered straight-through estimation (STE) theis2017lossy is applied for distortion computation, while additive uniform noise (AUN) balle2016end is used for entropy estimation during training.
 The module $g_{lrp}$ predicts quantization residuals based on decoded slices and hyperprior information:
@@
 Figure 3 illustrates the overall architecture of our proposed MLICv2 and MLICv2+.
-MLICv2 preserves the fundamental structure of MLIC++ while introducing several key enhancements to improve compression efficiency, particularly at high bit rates.
-First, we replace residual blocks following downsampling and upsampling operations with lightweight token mixing blocks, which enhance representational capacity while reducing computational complexity.
-Second, we upgrade the multi-reference entropy model to MEMv2 by incorporating enhanced inter-slice and intra-slice context modules with context reweighting and advanced positional embedding to better capture latent dependencies.
-Third, we introduce a selective compression module $f_{s}$ that intelligently predicts and skips zero elements in latent representations, reducing both bit rate and arithmetic coding time without compromising reconstruction quality.
-To demonstrate the potential for input-specific optimization, we propose MLICv2+, which iteratively refines latent features $\boldsymbol{y}$ and side information $\boldsymbol{z}$ based on the rate-distortion characteristics of individual inputs, enabling tailored compression strategies for different content types.
+MLICv2 retains the fundamental design of MLIC++ but introduces three key enhancements to improve compression efficiency, especially at high bit rates.
+First, residual blocks of transforms are replaced with lightweight token mixing blocks, enhancing representational capacity while reducing computational cost.
+Second, the multi-reference entropy model is upgraded to MEMv2, which integrates improved inter-slice and intra-slice context modules with context reweighting and advanced positional embeddings to better capture latent dependencies.
+Third, a selective compression module $f_{s}$ is introduced to predict and skip zero elements in latent representations, effectively lowering both bit rate and arithmetic coding time.
+Furthermore, MLICv2+ extends MLICv2 by incorporating iterative refinement of latent features $\boldsymbol{y}$ and side information $\boldsymbol{z}$ according to the rate-distortion characteristics of each input, enabling adaptive and content-specific compression strategies.
+3.3 Efficient Simple Token Mixing Transform (STMT)
+As illustrated in Figure 7, MLIC++ suffers from performance degradation at high bit rates due to the limited capacity of its transform modules to retain sufficient information in latent representations xu2022multi.
+The representational capacity is largely determined by the number of channels.
+Ballé et al.
+balle2020nonlinear showed that when channel numbers are small, the performance gap between low- and high-capacity models is negligible at low bit rates but becomes significant as the bit rate increases, with higher-capacity models achieving noticeably better reconstruction quality.
+Simply enlarging the channel dimension, however, leads to prohibitive computational costs.
+A more effective strategy is to enhance the transform design itself to improve capacity without substantially increasing complexity.
+Recent studies have explored transformer-based transforms liu2021swin, liu2023learned, li2024frequencyaware, zou2022the, lu2021tic, which achieve strong performance gains but often double the computational load compared to conventional residual blocks cheng2020learned, jiang2022mlic.
+Inspired by the MetaFormer framework yu2022metaformer, which separates token mixing and channel interaction, we design a lightweight token mixing block tailored for learned image compression.
+Unlike MetaFormer’s pooling-based token mixing that causes information loss, our design adopts a depth-wise residual block (DepthRB) jiang2024llic for nonlinear embedding, a depth-wise convolution for spatial mixing, and a gate block jiang2024llic for efficient channel interactions.
+Layer normalization (LN) ba2016layer is applied for training stability.
+Given an input feature $\boldsymbol{\omega}\in\mathbb{R}^{c\times h\times w}$, the process is formulated as:
+$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\mathrm{Conv}_{1\times 1}(\mathrm{DWConv}_{5\times 5}(\mathrm{DepthRB}(\mathrm{LN}(\boldsymbol{\omega})))),$ (6)
+$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\mathrm{Gate}(\mathrm{LN}(\boldsymbol{\omega})).$
+Stacking two token mixing blocks yields deeper transforms with enhanced expressiveness.
+Compared to conventional residual blocks with two $3\times 3$ convolutions cheng2020learned, our design achieves lower complexity and superior rate-distortion performance, particularly at high bit rates.
+Complexity analysis shows that stacking two token mixing blocks reduces MACs by 7.38% at $c=192$ and by 8.87% at $c=320$, compared to a single residual block cheng2020learned, jiang2023mlicpp.
 Figure 4: Visualization of the channels of hyperprior $\hat{\boldsymbol{H}}$ ($\lambda=0.013$).
 $\hat{\boldsymbol{H}}$ extracted from the latent representation of Image Kodim20 exhibits similar global similarity to original Kodim20.
-3.3 Efficient Simple Token Mixing Transform
-As illustrated in Figure 6, MLIC++ exhibits performance degradation at high bit rates due to limited capacity of transform modules to retain sufficient information in latent representations [76].
-The representational capacity is fundamentally influenced by the number of channels.
-As demonstrated by Ballé et al.
-[6], when channel numbers are limited, the performance gap between low- and high-capacity models is negligible at low bit rates.
-However, this gap becomes pronounced as bit rates increase, with higher-capacity models achieving significantly better reconstruction quality.
-Simply increasing channel numbers often results in prohibitive computational complexity.
-A more practical approach is to enhance the architectural design itself to improve transform capacity without substantially increasing computational cost.
-Recent work has adopted transformer layers [46] for transform coding [45, 40, 88, 48], demonstrating remarkable performance improvements.
-However, existing techniques [45] can introduce up to twice the complexity of original residual blocks in the MLIC series.
-Inspired by the elegance of MetaFormer [81], which consists of token mixing and channel-wise interaction components, we investigate its effectiveness for LIC.
-However, MetaFormer’s use of pooling for token mixing, while efficient, is suboptimal for compression tasks due to information loss.
-We propose a simple token mixing block as illustrated in Figure 2.
-For token mixing, we employ a depth-wise residual block (DepthRB) [32] for nonlinear embedding, depth-wise convolution for spatial token mixing, and a gate block [32] for efficient channel-wise interactions.
-LayerNorm (LN) [5] is adopted for stable training.
-Given input feature $\boldsymbol{\omega}$, the process is formulated as:
-$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\textrm{Conv}1\times 1(\textrm{DWConv}5\times 5(\textrm{DepthRB}(\textrm{LN}(\boldsymbol{\omega})))),$ (7)
-$\boldsymbol{\omega}$ $=\boldsymbol{\omega}+\textrm{Gate}(\textrm{LN}(\boldsymbol{\omega})),$
-where $\boldsymbol{\omega}\in\mathbb{R}^{c\times h\times w}$ represents the input feature with $c$, $h$, $w$ denoting channels, height, and width respectively.
-Stacking two token mixing blocks creates deeper transforms with enhanced expressiveness.
-Compared to previous residual blocks [16] consisting of two $3\times 3$ convolutions, our token mixing block achieves lower complexity while delivering significantly better performance, especially at high bit rates.
-The complexity analysis reveals that stacking two simple token-mixing blocks reduces multiply-accumulate operations (MACs) by 7.38% at $c=192$ and by 8.87% at $c=320$, compared with stacking a single residual block [16, 34, 33].
 3.4 Enhanced Multi-Reference Entropy Modeling
-The enhanced multi-reference entropy model (MEMv2) is presented in Figure 2.
-To address the limitations of MLIC++, we introduce several key innovations: hyperprior-based global correlation prediction, context reweighting to handle varying channel importance, 2D Rotary Positional Embedding (RoPE) and guided selective compression for improved context modeling and entropy coding.
-3.4.1 Hyperprior-Guided Global Correlation Prediction
-In MLIC++ [33], global similarity between anchor and non-anchor parts of the previous slice is exploited to predict current slice similarities.
+The enhanced multi-reference entropy model (MEMv2) is illustrated in Figure 3.
+To overcome the limitations of MLIC++, MEMv2 incorporates several key innovations: (i) hyperprior-guided global correlation prediction (HGCP) for more accurate dependency modeling, (ii) context reweighting (CR) to adaptively emphasize informative channels, (iii) 2D Rotary Positional Embedding (2D RoPE) for enhanced spatial awareness, and (iv) guided selective compression (GSC) to improve entropy coding efficiency.
+Together, these components enable more expressive and adaptive context modeling, particularly at high bit rates.
+3.4.1 Hyperprior-Guided Global Correlation Prediction (HGCP)
+In MLIC++ jiang2023mlicpp, global similarity between anchor and non-anchor parts of the previous slice is exploited to predict current slice similarities.
 However, for the first slice $\hat{\boldsymbol{y}}^{0}$, no such reference exists.
@@
 We leverage this property by estimating global similarity between $\hat{\boldsymbol{y}}_{ac}^{0}$ and $\hat{\boldsymbol{y}}_{na}^{0}$ using the similarity between anchor and non-anchor parts of the hyperprior, $\hat{\boldsymbol{H}}_{ac}$ and $\hat{\boldsymbol{H}}_{na}$.
-Following MLIC++, we employ linear attention [60] for global context modeling and introduce a Gate block [32] for enhanced pointwise interaction:
+Following MLIC++, we employ linear attention shen2021efficient for global context modeling and employ a Gate block introduced in LLIC jiang2024llic for enhanced pointwise interaction:
 $\hat{\boldsymbol{C}}_{g}^{0}$ $=\textrm{Softmax}_{2}(\hat{\boldsymbol{H}}_{na,q})\left(\textrm{Softmax}_{1}(\hat{\boldsymbol{H}}_{ac,k})^{\top}\hat{\boldsymbol{y}}^{0}_{g,v}\right),$ (7)
 $\hat{\boldsymbol{C}}_{g}^{0}$ $=\textrm{Gate}(\hat{\boldsymbol{C}}_{g}^{0})+\hat{\boldsymbol{C}}_{g}^{0},$
-where $\hat{\boldsymbol{H}}_{na,q}=\text{Linear}(\hat{\boldsymbol{H}}_{na})$, $\hat{\boldsymbol{H}}_{ac,k}=\text{Linear}(\hat{\boldsymbol{H}}_{ac})$, and $\hat{\boldsymbol{y}}^{0}_{g,v}=\text{Linear}(\hat{\boldsymbol{y}}_{ac}^{0})$.
-While prior methods [30, 34, 33, 32] capture only local dependencies in the first slice, our hyperprior-guided mechanism models global dependencies effectively.
-Given the information compaction property of latent representations [30], the fact that the first slice typically exhibits higher entropy as it serves as reference for subsequent slices, incorporating global context significantly reduces bit rate and improves compression performance.
+where $\hat{\boldsymbol{H}}_{na,q}={\text{LinearLayer}}(\hat{\boldsymbol{H}}_{na})$, $\hat{\boldsymbol{H}}_{ac,k}={\text{LinearLayer}}(\hat{\boldsymbol{H}}_{ac})$, and $\hat{\boldsymbol{y}}^{0}_{g,v}={\text{LinearLayer}}(\hat{\boldsymbol{y}}_{ac}^{0})$.
+While prior methods he2022elic, jiang2022mlic, jiang2023mlicpp, jiang2024llic capture only local dependencies in the first slice, our hyperprior-guided mechanism effectively models global dependencies.
+Due to the information compaction property of latent representations he2022elic, the first slice typically exhibits higher entropy as it serves as a reference for subsequent slices.
+Incorporating global context at this stage thus significantly reduces bit rate and enhances compression performance.
 Figure 5: Proposed context reweighting module.
-3.4.2 Context Reweighting
-In MLIC++ [33], the contexts are extracted by spatial attention mechanism.
-Given the query $\boldsymbol{Q}\in\mathbb{R}^{L\times c}$, key $\boldsymbol{K}\in\mathbb{R}^{L\times c}$ and the value $\boldsymbol{V}\in\mathbb{R}^{L\times c}$, where $L=hw$ is the sequence length, $h$ is the height, $w$ is the width, and $c$ is the channel number, in most spatial attention [70, 60], the attention map is computed by $\textrm{Sim}(\boldsymbol{Q}\boldsymbol{K}^{\top})\in\mathbb{R}^{L\times L}$, where Sim is the predefined function, such as Softmax [70] or two independent Softmax operations [60], indicating the information is gathered and weighted from spatial tokens ($\in\mathbb{R}^{1\times c}$).
+3.4.2 Context Reweighting (CR)
+In MLIC++ jiang2023mlicpp, the contexts are extracted by spatial attention mechanism.
+Given the query $\boldsymbol{Q}\in\mathbb{R}^{L\times c}$, key $\boldsymbol{K}\in\mathbb{R}^{L\times c}$ and the value $\boldsymbol{V}\in\mathbb{R}^{L\times c}$, where $L=hw$ is the sequence length, $h$ is the height, $w$ is the width, and $c$ is the channel number, in most spatial attention vas2017attention, shen2021efficient, the attention map is computed by $\textrm{Sim}(\boldsymbol{Q}\boldsymbol{K}^{\top})\in\mathbb{R}^{L\times L}$, where Sim is the predefined function, such as Softmax vas2017attention or two independent Softmax operations shen2021efficient, indicating the information is gathered and weighted from spatial tokens ($\in\mathbb{R}^{1\times c}$).
 However, one common limitation is that all channel-wise features are treated equally, without considering that different channels may carry varying levels of importance.
 We argue that adaptively reweighting channel-wise context features can lead to more accurate entropy modeling.
-By suppressing noisy or less informative channels and emphasizing more relevant ones, the model has the potential to capture context information more precisely and improve overall compression performance, which has not been considered in existing methods [34, 33, 30, 31, 51, 19, 22, 54, 55].
-To address this issue, we introduce a two-stage context modeling strategy, where spatial aggregation [33] is first performed, followed by adaptive channel-wise context reweighting to capture the varying importance of different features.
+By suppressing noisy or less informative channels and emphasizing more relevant ones, the model has the potential to capture context information more precisely and improve overall compression performance, which has not been considered in existing methods jiang2022mlic, jiang2023mlicpp, he2022elic, he2021checkerboard, minnen2020channel, duan2023qarv, fu2023learned, qian2022entroformer, qian2020learning.
+To address this issue, we introduce a two-stage context modeling strategy, where spatial aggregation jiang2023mlicpp is first performed, followed by adaptive channel-wise context reweighting to capture the varying importance of different features.
 The architecture is in Fig.
-4.
-Specifically, in context reweighting, the context is transformed to query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$, and the query and values are transposed to compute the channel-wise attention map [3].
+5.
+Specifically, in context reweighting, the context is transformed to query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$, and the query and values are transposed to compute the channel-wise attention map ali2021xcit.
 The process is formulated as:
-$\boldsymbol{O}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})\boldsymbol{V}^{\top}.$
+$\boldsymbol{O}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})\boldsymbol{V}^{\top}.$ (8)
+The attention map $\boldsymbol{M}=\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})$ is row-wise normalized via the Softmax function: $\boldsymbol{M}_{i,j}=\frac{\exp(\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j})}{\sum_{j^{\ast}=1}^{C}\exp(\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j^{\ast}})},$ where $\boldsymbol{q}_{i}$ and $\boldsymbol{k}_{j}$ are the $i$-th and $j$-th channel vectors of $\boldsymbol{Q}$ and $\boldsymbol{K}$, respectively.
+The summation is over spatial indices while channel indices remain open.
 The resolution of the attention map $\textrm{Softmax}(\boldsymbol{Q}^{\top}\boldsymbol{K})$ is $c\times c$, indicating the importance of each channel ($\in\mathbb{R}^{L\times 1}$).
@@
 (8) is $\mathcal{O}(c^{2}hw)$.
-$\boldsymbol{O}$ is then fed into a Gate block [32] for efficient point-wise interactions.
+$\boldsymbol{O}$ is then fed into a Gate block jiang2024llic for efficient point-wise interactions.
 This attention-based reweighting helps the model focus on the most relevant parts of the context.
@@
 Our approach brings consistent performance improvements with minimal additional cost.
-3.4.3 2D Rotary Position Embedding
-MLIC++ [33] applies relative position embedding (RPE) from Swin-Transformer [46] directly to attention maps as additional bias.
+3.4.3 2D Rotary Position Embedding (2D RoPE)
+MLIC++ jiang2023mlicpp applies relative position embedding (RPE) from Swin-Transformer liu2021swin directly to attention maps as additional bias.
 However, this direct addition may limit interaction with attention weights.
-Recently, Rotary Positional Embedding (RoPE) [63], a multiplication-based approach, has been widely adopted in language models [66, 58].
+Recently, Rotary Positional Embedding (RoPE) su2024roformer, a multiplication-based approach, has been widely adopted in language models touvron2023llama, reid2024gemini.
 RoPE offers several advantages: compatibility with linear attention (whereas RPE requires adding biases that conflict with implicit attention computation), better extrapolation capabilities, and more concise implementation.
-For 2D image content, we extend the original 1D RoPE to handle horizontal, vertical, and diagonal relationships.
+For 2D image content, we extend the original 1D RoPE to handle horizontal (width), vertical (height), and diagonal relationships.
 Given query $\boldsymbol{q}_{m_{x},m_{y}}$ and key $\boldsymbol{k}_{n_{x},n_{y}}$ at 2D positions $(m_{x},m_{y})$ and $(n_{x},n_{y})$, we rotate queries and keys twice with different angles $\theta_{x}$ and $\theta_{y}$ for horizontal and vertical directions.
@@
 where $\theta_{(m_{x}-n_{x},m_{y}-n_{y})}=(m_{x}-n_{x})\theta_{x}+(m_{y}-n_{y})\theta_{y}$ encodes relative positions.
-Due to the ambiguity in defining relative importance of horizontal versus vertical positional dependencies, we make angles $\theta_{x}$ and $\theta_{y}$ learnable, initializing them to 10000 following original RoPE [63].
-3.4.4 Guided Selective Compression
+Due to the ambiguity in defining relative importance of horizontal versus vertical positional dependencies, we make angles $\theta_{x}$ and $\theta_{y}$ learnable, initializing them to 10000 following original RoPE su2024roformer.
+For $c\textgreater 2$, the queries and keys are divided into $\frac{c}{2}$ groups, each following Equation (10).
+3.4.4 Guided Selective Compression (GSC)
 Zero-centered quantization in Equation 4 produces many zero elements in $\lceil\boldsymbol{\Lambda}\rfloor=\lceil\boldsymbol{y}-\boldsymbol{\mu}\rfloor$, as illustrated in Figure 6.
 While these zeros consume minimal bit rate, they account for significant arithmetic encoding/decoding time.
-Recent selective compression methods [39, 2] have employed neural networks to predict selective compression maps, but they are primarily applied to simpler entropy models.
-In addition, scale-based prediction has been introduced in recent works [62, 84].
-However, when extended to advanced entropy models [34, 33], the training phase becomes highly unstable, with the overall loss value becoming “nan” after several steps, regardless of whether the parameters of the compression model are frozen or not.
+Recent selective compression methods lee2022selective, alexandre2023hierarchical have employed neural networks to predict selective compression maps, but they are primarily applied to simpler entropy models.
+In addition, scale-based prediction has been introduced in recent works shi2022alphavc, zhang2024practical.
+However, when extended to advanced entropy models jiang2022mlic, jiang2023mlicpp, the training phase becomes highly unstable, with the overall loss value becoming “nan” after several steps, regardless of whether the parameters of the compression model are frozen or not.
 The observed training instability is primarily caused by the dependency introduced by the context modeling, rather than by the selective compression mechanism itself.
 Specifically, skipping non-zero elements in the current slice can significantly affect the entropy modeling of subsequent slices due to the inter-latent dependencies established by the context model.
-Considering the substantial performance gains enabled by context modeling, our goal is to skip the encoding of zero elements as much as possible—on top of the existing context modeling framework [51, 30, 33, 34]—in order to reduce the time spent on arithmetic encoding and decoding.
+Considering the substantial performance gains enabled by context modeling, our goal is to skip the encoding of zero elements as much as possible—on top of the existing context modeling framework minnen2020channel, he2022elic, jiang2023mlicpp, jiang2022mlic—in order to reduce the time spent on arithmetic encoding and decoding.
 To address above issues, the guided selective compression module $f_{s}$ is proposed for post-training.
@@
 The architecture of the guided selective compression $f_{s}$ module is presented in Fig.
-2.
+3.
 Specifically, in $f_{s}$, the scale-based selective prediction is employed as the initial predicted selective compression map.
-For $j$-th element ${s}_{\boldsymbol{\sigma},j}$ in $\boldsymbol{s}_{\boldsymbol{\sigma}}$, the scale-based selective compression value is:
+For $j$-th element ${s}_{\boldsymbol{\sigma},j}$ in $\boldsymbol{s}_{\boldsymbol{\sigma}}$, the scale-based selective compression value is served as the initialization:
 ${s}_{\boldsymbol{\sigma},j}=\left\{\begin{aligned} &0,\quad{\sigma}_{j}\textless{\xi},\\ &1,\quad{\sigma}_{j}\geq{\xi},\end{aligned}\right$. (11)
@@
 The hyperprior $\hat{\boldsymbol{H}}$ also serves as a prior.
-The predicted selective map is formulated as:
-$\boldsymbol{\varepsilon}^{i}_{ac}$ $=f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},ac}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\hat{\boldsymbol{H}}\right),\boldsymbol{s}^{i}_{ac}=\textrm{STE}\left(\textrm{Sigmoid}\left(\boldsymbol{\varepsilon}^{i}_{ac}\right)\right),$ (12)
-$\boldsymbol{\varepsilon}^{i}_{na}$ $=f_{s}\Bigl{(}\boldsymbol{s}_{\boldsymbol{\sigma},na}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\boldsymbol{s}^{i}_{ac},\hat{\boldsymbol{H}}\Bigr{)},\boldsymbol{s}^{i}_{na}=\textrm{STE}\left(\textrm{Sigmoid}\left(\boldsymbol{\varepsilon}^{i}_{na}\right)\right).$
+The scale-based initialization is refined by $f_{s}$ for final skip map:
+$\boldsymbol{\varepsilon}^{i}_{ac}$ $={\textrm{Sigmoid}}\left(f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},ac}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\hat{\boldsymbol{H}}\right)\right),\boldsymbol{s}^{i}_{ac}=\textrm{STE}\left(\boldsymbol{\varepsilon}^{i}_{ac}\right),$ (12)
+$\boldsymbol{\varepsilon}^{i}_{na}$ $={\textrm{Sigmoid}}\left(f_{s}\left(\boldsymbol{s}_{\boldsymbol{\sigma},na}^{i},\lceil\boldsymbol{\Lambda}^{\textless i}\rfloor\odot\boldsymbol{s}^{\textless i},\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\boldsymbol{s}^{i}_{ac},\hat{\boldsymbol{H}}\right)\right),\boldsymbol{s}^{i}_{na}=\textrm{STE}\left(\boldsymbol{\varepsilon}^{i}_{na}\right).$
 During training, $\lceil\boldsymbol{\Lambda}\rfloor$ is known, therefore cross-entropy loss is employed to optimize $f_{s}$, treating selective prediction as a binary classification task.
-Additionally, elements with larger values have a higher impact on distortion, so the cross-entropy loss is weighted by $\lceil\boldsymbol{\Lambda}\rfloor+1$.
-To the best of our knowledge, we are the first to highlight the importance of quantization residuals and incorporate a weighted loss accordingly.
+Additionally, elements with larger values have a higher impact on distortion, so the cross-entropy loss is weighted by $|\lceil\boldsymbol{\Lambda}\rfloor|+1$.
 The overall loss is:
@@
 There are many zero elements in $\lceil\boldsymbol{\Lambda}\rfloor$.
-$\mathcal{L}_{s}=\sum_{i=0}^{L}\Bigl{(}(\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor+1)\odot\mathcal{L}^{i}_{ce,ac}+(\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor+1)\odot\mathcal{L}^{i}_{ce,na}\Bigr{)},$ (13)
-$\mathcal{L}_{ce,ac}^{i}=-\frac{\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\odot\log\boldsymbol{\varepsilon}^{i}_{ac}+\left(1-\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor\right)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{ac})}{2},$
-$\mathcal{L}_{ce,na}^{i}=-\frac{\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor\odot\log\boldsymbol{\varepsilon}^{i}_{na}+\left(1-\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor\right)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{na})}{2}.$
-3.5 Iterative Latent Refinement
+${\mathcal{L}_{s}=\sum_{i=0}^{L}\Bigl((|\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor|+1)\odot\mathcal{L}^{i}_{ce,ac}+(|\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor|+1)\odot\mathcal{L}^{i}_{ce,na}\Bigr)},$ (13)
+${\mathcal{L}_{ce,ac}^{i}=\tfrac{1}{2}\!\left[-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor|>0\big)\odot\log\boldsymbol{\varepsilon}^{i}_{ac}-\big(1-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{ac}\rfloor|>0\big)\big)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{ac})\right]},$
+${\mathcal{L}_{ce,na}^{i}=\tfrac{1}{2}\!\left[-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor|>0\big)\odot\log\boldsymbol{\varepsilon}^{i}_{na}-\big(1-\mathbf{1}\!\big(|\lceil\boldsymbol{\Lambda}^{i}_{na}\rfloor|>0\big)\big)\odot\log(1-\boldsymbol{\varepsilon}^{i}_{na})\right]}.$
+Since both the encoder and decoder compute identical selective compression maps using $f_{s}$ with the same inputs, the selective coding decisions are perfectly synchronized without requiring any additional information transmission or bitstream overhead.
+To the best of our knowledge, we are the first to highlight the importance of quantization residuals and incorporate a weighted loss accordingly.
+3.5 Iterative Latent Refinement (ILR)
 To validate our MLICv2’s potential and demonstrate achievable performance improvements, we employs encoding-time scaling to find optimal quantized latent representation $\hat{\boldsymbol{y}}$ and side information $\hat{\boldsymbol{z}}$ for each input.
 Direct encoder optimization requires storing model weights, gradients and intermediate features, leading to prohibitive memory consumption for high-resolution images.
-We adopt Stochastic Gumbel Annealing (SGA) [80] to refine $\boldsymbol{y}$ and $\boldsymbol{z}$ efficiently.
+We adopt Stochastic Gumbel Annealing (SGA) yang2020improving to refine $\boldsymbol{y}$ and $\boldsymbol{z}$ efficiently.
 In SGA, outputs of $g_{a}$ and $h_{a}$ serve as initialization.
@@
 we primarily employ it as a validation tool to demonstrate our model’s potential for instance-level adaptability and performance improvements.
-Rate control mechanisms [84] could be incorporated in future work for applications requiring strict bit-rate constraints.
+Rate control mechanisms zhang2024practical could be incorporated in future work for applications requiring strict bit-rate constraints.
 Figure 7: PSNR-Bit-rate curve and MS-SSIM-Bit-rate curve. MS-SSIM is converted to dB for better visual clarity. Please zoom in for better view.
+Table 1: BD-Rate $(\%)$ comparison for PSNR (dB) and MS-SSIM. The anchor is VTM-17.0 Intra. “-” means the result is not available.
 Methods Venue BD-Rate (%) w.r.t. VTM 17.0 Intra
-   Kodak    Tecnick    CLIC Pro Valid
-   PSNR    MS-SSIM    PSNR    PSNR
-   Cheng’20 [16]    CVPR’20    5.58    –44.21    7.57    11.71
-   ChARM [51]    ICIP’20    3.23    –    –0.88    –
-   Qian’21 [55]    ICLR’21    10.05    –39.53    7.52    0.02
-   Xie’21 [75]    ACMMM’21    1.55    –43.39    -0.80    3.21
-   NLAIC’21 [15]    TIP’21    11.80    –40.27    7.11    –
-   Gao’21 [24]    ICCV’21    –8.76    –48.60    -    –6.49
-   Guo’22 [26]    TCSVT’22    –4.45    –45.23    –    –
-   LBHIC [74]    TCSVT’22    –4.56    –50.54    –    –
-   Entroformer [54]    ICLR’22    4.73    –42.64    2.31    –1.04
-   SwinT-ChARM [87]    ICLR’22    –1.73    –    –    –
-   NeuralSyntax [71]    CVPR’22    8.97    –39.56    –    5.64
-   STF [88]    CVPR’22    –2.48    –47.72    –2.75    0.42
-   WACNN [88]    CVPR’22    –2.95    –47.71    –5.09    0.04
-   ELIC [30]    CVPR’22    –5.95    –44.60    –9.14    –3.45
-   Contextformer [38]    ECCV’22    –5.77    –46.12    –9.05    –
-   GLLMM [22]    TIP’23    –1.43    –47.41    –5.58    –
-   NVTC [21]    CVPR’23    –1.04    –    –    –
-   LIC-TCM Large [45]    CVPR’23    –10.14    –48.94    –11.47    –8.04
-   QARV [19]    TPAMI’24    0.31    –    –3.03    -
-   FTIC [40]    ICLR’24    –12.99    –51.13    –14.88    –9.53
-   LLIC-TCM [32]    TMM’24    –10.94    –49.73    –14.99    –10.41
-   WeConvene [23]    ECCV’24    –6.71    –48.63    –8.42    –5.74
-   CCA [27]    NeurIPS’24    –12.04    –    –14.55    –10.75
-   MambaIC [82]    CVPR’25    –13.03    –    –18.41    –15.58
-   LALIC [20]    CVPR’25    –13.72    –50.86    –18.19    –14.34
-   DCAE [47]    CVPR’25    –15.43    –53.20    –20.51    –15.78
-   Baseline Models
-   MLIC [34]    ACMMM’23    -8.05    –49.13    –12.73    –8.79
-   MLIC+ [34]    ACMMM’23    -11.39    –52.75    –16.38    –12.56
-   MLIC++ [33]    ICMLW’23    -13.39    -53.63    -17.59    -13.08
-   MLICv2    Ours    -16.54    -54.56    -21.61    -16.05
-   MLICv2+    Ours    -20.46    -55.86    -24.35    -19.14
-Table 1: BD-Rate $(\%)$ comparison for PSNR (dB) and MS-SSIM. The anchor is VTM-17.0 Intra. “-” means the result is not available.
-Figure 7: Reconstructions of our proposed MLICv2, recent learned image compression models [33, 30, 88, 16] and VTM-17.0 Intra [12].
-“[MSE]” denotes the model is optimized for MSE.
-Please zoom in for better view.
-(1) In MLICv2+∗, gradient checkpointing [14] is employed during the SGA iterations.
-(2) The MACs of MLICv2+ are theoretically estimated based on the number of SGA iterations, following the observation from Epoch AI that backward and forward passes almost always have MAC ratios close to 2:122footnotemark: 2.
-The MACs of MLICv2+∗ are not reported due to the use of gradient checkpointing, which prevents accurate profiling.
-(3) The lack of reported results for some models at certain resolutions is due to out-of-memory (OOM) issues or inaccurate profiling.
-Figure 8: GPU memory consumption, Forward MACs, encoding time and decoding time comparisons among our proposed models and recent learned image compression models  [16, 75, 71, 88, 30, 45, 40, 34, 33].
+Kodak kodak Tecnick tecnick2014TESTIMAGES CLIC Pro Valid CLIC2020
+PSNR MS-SSIM PSNR MS-SSIM PSNR MS-SSIM
+MLIC jiang2022mlic ACMMM’23 -8.05 -49.13 -12.73 -47.26 -8.79 -45.79
+MLIC+ jiang2022mlic ACMMM’23 -11.39 -52.75 -16.38 -53.54 -12.56 -48.75
+MLIC++ jiang2023mlicpp ICMLW’23 -13.39 -53.63 -17.59 -53.83 -13.08 -50.78
+MLICv2 Ours -16.54 -54.56 -21.61 -55.26 -16.05 -51.86
+MLICv2+ Ours -20.46 -55.86 -24.35 -56.42 -19.14 -52.75
 4 Experiments
@@
 Our training dataset contains 101,104 images with resolutions exceeding $640\times 640$ pixels.
-These images are carefully selected from various established image datasets [18, 43, 1, 42, 77].
+These images are carefully selected from various established image datasets deng2009imagenet, lin2014microsoft, Agustsson2017NTIRE2C, lim2017enhanced, yang2023hq.
 To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al.
-[10] by further down-sampling the JPEG images using a randomized factor.
+balle2018variational by further down-sampling the JPEG images using a randomized factor.
 Additionally, images with bits per pixel (bpp) lower than 3 are excluded to ensure data quality.
@@
 MLICv2 is implemented using PyTorch 2.2.2 and CompressAI 1.2.6.
-Following CompressAI, we configure $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE optimization and $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) [73].
+Following CompressAI, we configure $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE optimization and $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) wang2003multiscale.
 Training is conducted on eight Tesla A100-80G GPUs with a batch size of 16, employing the Adam optimizer $\beta_{1}=0.9,\beta_{2}=0.999$.
@@
 4.2 Benchmarks and Metrics
-We evaluate rate-distortion performance on three datasets extensively used in learned compression research [9, 64, 10, 50, 51, 16, 75, 31, 30, 88, 87, 38, 19, 71, 26, 74, 15, 45, 34, 33, 32]:
-  • Kodak [37]: Contains 24 uncompressed images with $768\times 512$ pixels.
-  • Tecnick [4]: Comprises 100 uncompressed images with $1200\times 1200$ pixels.
-  • CLIC Pro Valid [65]: The validation set from the 3rd Challenge on Learned Image Compression, consisting of 41 high-resolution images with $2048\times 1440$ pixels.
-Performance ranking employs the Bjøntegaard delta rate (BD-Rate) metric [11] for comprehensive evaluation.
+We evaluate rate-distortion performance on three datasets extensively used in learned compression research balle2016end, theis2017lossy, balle2018variational, minnen2018joint, minnen2020channel, cheng2020learned, xie2021enhanced, he2021checkerboard, he2022elic, zou2022the, zhu2022transformerbased, koyuncu2022contextformer, duan2023qarv, wang2022neural, guo2021causal, wu22block, chen2021nic, liu2023learned, jiang2022mlic, jiang2023mlicpp, jiang2024llic:
+  • Kodak kodak: Contains 24 uncompressed images with $768\times 512$ pixels.
+  • Tecnick tecnick2014TESTIMAGES: Comprises 100 uncompressed images with $1200\times 1200$ pixels.
+  • CLIC Pro Valid CLIC2020: The validation set from the 3rd Challenge on Learned Image Compression, consisting of 41 high-resolution images with $2048\times 1440$ pixels.
+Performance ranking employs the Bjøntegaard delta rate (BD-Rate) metric bjontegaard2001calculation for comprehensive evaluation.
+Figure 8: Reconstructions of our proposed MLICv2, recent learned image compression models jiang2023mlicpp, he2022elic, zou2022the, cheng2020learned, feng2025linear, jiang2024llic and VTM-17.0 Intra bross2021vvc.
+“[MSE]” denotes the model is optimized for MSE.
+Please zoom in for better view.
+(1) In MLICv2+∗, gradient checkpointing chen2016training is employed during the SGA iterations.
+(2) The MACs of MLICv2+ are theoretically estimated based on the number of SGA iterations, following the observation from Epoch AI that backward and forward passes almost always have MAC ratios close to 2:1444https://epoch.ai/blog/backward-forward-FLOP-ratio.
+The MACs of MLICv2+∗ are not reported due to the use of gradient checkpointing, which prevents accurate profiling.
+(3) The lack of reported results for some models at certain resolutions is due to out-of-memory (OOM) issues or inaccurate profiling.
+Figure 9: GPU memory consumption, Forward kMACs/pixel, encoding time and decoding time comparisons among our proposed models and recent learned image compression models  cheng2020learned, xie2021enhanced, wang2022neural, zou2022the, he2022elic, liu2023learned, li2024frequencyaware, jiang2022mlic, jiang2023mlicpp.
 4.3 Rate-Distortion Performance
@@
 For a fair comparison with learned image compression models, VTM-17.0 Intra is evaluated in YUV444 color space under encoder_intra_vtm.cfg.
-To demonstrate the superiority of our MLICv2 and MLICv2+, we compare with recent learned image compression models [41, 16, 74, 51, 55, 75, 15, 26, 54, 87, 88, 30, 38, 22, 21, 45, 40, 32, 23, 20, 47].
+To demonstrate the superiority of our MLICv2 and MLICv2+, we compare with recent learned image compression models li2020efficient, cheng2020learned, wu22block, minnen2020channel, qian2020learning, xie2021enhanced, chen2021nic, guo2021causal, qian2022entroformer, zhu2022transformerbased, zou2022the, he2022elic, koyuncu2022contextformer, fu2023learned, feng2023nvtc, liu2023learned, li2024frequencyaware, jiang2024llic, fu2024weconvene, feng2025linear, lu2025learned.
 The rate-distortion curves are presented in Fig.
-6.
-When compared with our baseline MLIC++, MLICv2+ performs significantly better, especially at high bit-rates and on high-resolution images [4, 65].
+7.
+When compared with our baseline MLIC++, MLICv2+ performs significantly better, especially at high bit-rates and on high-resolution images tecnick2014TESTIMAGES, CLIC2020.
 Specifically, our MLICv2 achieves a maximum improvement of $0.6,0.5,0.5$ dB when the bpp is around $0.8,0.55,0.6$ on Kodak, Tecnick, CLIC Pro Val, respectively.
-Our MLICv2+ achieves larger improvement compared to MLIC++ [33] due to adopted latent refinement [80] for enhanced instance adaptability.
-The average improvements of latent refinement on MLICv2 on Kodak [37], Tecnick [4], CLIC Pro Val [65] are $0.22,0.15,0.15$ dB, respectively.
+Our MLICv2+ achieves larger improvement compared to MLIC++ jiang2023mlicpp due to adopted latent refinement yang2020improving for enhanced instance adaptability.
+The average improvements of latent refinement on MLICv2 on Kodak kodak, Tecnick tecnick2014TESTIMAGES, CLIC Pro Val CLIC2020 are $0.22,0.15,0.15$ dB, respectively.
 BD-Rate reductions over VTM-17.0 Intra are presented in Table 1.
 Compared with existing methods, our MLICv2 and MLICv2+ achieve state-of-the-art performance, reducing BD-Rates by $16.54\%$ and $20.46\%$ on Kodak.
-When compared to the baseline model MLIC++ [33], our MLICv2 and MLICv2+ reduce $3.15\%$ and $7.07\%$ more bit-rates over VTM-17.0 Intra on Kodak [37], respectively.
-Our MLICv2 and MLICv2+ also achieve superior performances on high-resolution datasets Tecnick [4] and CLIC Pro Val [65].
-Specifically, our MLICv2 and MLICv2+ reduce BD-Rates by $21.61\%$ and $24.35\%$ on Tecnick [4] and $16.05\%$ and $19.14\%$ on CLIC Pro Val [65].
+When compared to the baseline model MLIC++ jiang2023mlicpp, our MLICv2 and MLICv2+ reduce $3.15\%$ and $7.07\%$ more bit-rates over VTM-17.0 Intra on Kodak kodak, respectively.
+Our MLICv2 and MLICv2+ also achieve superior performances on high-resolution datasets Tecnick tecnick2014TESTIMAGES and CLIC Pro Val CLIC2020.
+Specifically, our MLICv2 and MLICv2+ reduce BD-Rates by $21.61\%$ and $24.35\%$ on Tecnick tecnick2014TESTIMAGES and $16.05\%$ and $19.14\%$ on CLIC Pro Val CLIC2020.
 4.3.2 Qualitative Results
-We also compare the qualitative performance to MLIC++, ELIC [30], STF, WACNN [88], Cheng’20 [16] and VTM-17.0 Intra on subjective qualities.
+We also compare the qualitative performance to MLIC++, ELIC he2022elic, STF, WACNN zou2022the, Cheng’20 cheng2020learned and VTM-17.0 Intra on subjective qualities.
 The windowsill of reconstructions are cropped to patches for clearer comparisons.
@@
 In terms of visual quality, our MLICv2 has significant improvements on rate-perception performance compared to other models.
-4.4 Computational Complexity
-The model complexities and sizes are presented in Fig.
-8 and Table 5, respectively.
-To comprehensively evaluate model complexity, 16 images from the LIU4K dataset [44] are cropped to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072\}$ patches to cover the various resolutions of images that can be encountered in practical scenarios.
-Forward GPU memory, encoding time, decoding time, and forward Multiply-Accumulate operations (MACs) are employed as the complexity metrics.
-Encoding and decoding times include arithmetic encoding and decoding times.
-Since the number of skipped elements at different bit-rates varies, we report the results of MLICv2 and MLICv2+ with $\lambda=0.0130$.
-Due to the high memory consumption introduced by SGA, we employ gradient checkpointing [14] to overcome GPU memory limitations.
-This technique alleviates memory pressure by selectively recomputing a subset of intermediate activations during the backward pass.
-Instead of storing all intermediate activations during the forward pass, only a small set—typically those essential for gradient computation—are cached.
-The remaining activations are recomputed on-the-fly during backpropagation.
-This strategy significantly reduces peak memory usage, albeit at the cost of increased computation time.
 (a) Kodim01
@@
 The attention map is computed by $\textrm{Softmax}_{2}(\hat{\boldsymbol{H}}_{na,q})[j]\textrm{Softmax}_{1}(\hat{\boldsymbol{H}}_{ac,k})^{\top}$, where $j$ is the query index.
-11footnotetext: https://epoch.ai/blog/backward-forward-FLOP-ratio
+4.4 Computational Complexity
+The model complexities and sizes are summarized in Fig.
+9 and Table 5, respectively.
+To evaluate complexity across resolutions, 16 images from the LIU4K dataset liu2020comprehensive are cropped to patches of $\{512^{2},768^{2},1024^{2},1536^{2},2048^{2},2560^{2},3072^{2}\}$ pixels, covering a wide range of practical image sizes.
+We report forward GPU memory, encoding and decoding times (including arithmetic coding), and forward kMACs/pixel as complexity metrics.
+Since the number of skipped elements varies with bit rate, results for MLICv2 and MLICv2+ are reported at $\lambda=0.0130$.
+Due to the high memory consumption of SGA, we adopt gradient checkpointing chen2016training to alleviate GPU memory limitations.
+Instead of storing all intermediate activations during the forward pass, only a small subset necessary for gradient computation is cached, while the remaining activations are recomputed on-the-fly during backpropagation.
+This strategy substantially reduces peak memory usage at the cost of increased computation time.
 4.4.1 On Forward GPU Memory
-Compared to MLIC+[34], MLICv2 significantly reduces memory usage thanks to its linear-complexity entropy model.
-Relative to MLIC++[33], it incurs slightly higher memory for high-resolution images due to the additional modules introduced in this work, but still requires less memory than Xie’21 [75], Qian’21 [55], Entroformer [55], and LIC-TCM [45].
+Compared to MLIC+jiang2022mlic, MLICv2 significantly reduces memory usage thanks to its linear-complexity entropy model.
+Relative to MLIC++jiang2023mlicpp, it incurs slightly higher memory for high-resolution images due to the additional modules introduced in this work, but still requires less memory than Xie’21 xie2021enhanced, Qian’21 qian2020learning, Entroformer qian2020learning, and LIC-TCM liu2023learned.
 This modest increase is justified by the corresponding performance improvements.
@@
 Figure 11: Upper-bound reconstruction quality (PSNR) on the Kodak dataset without entropy constraints.
-Transform Architecture Cheng’20 [16] SwinT [87] ELIC [30] MLIC++ LIC-TCM [45] MLICv2
-BD-Rate(%) 0.00 -4.61 -5.19 -0.41 -8.91 -9.44
-Table 2: Comparisons among different transforms of proposed MLICv2, LIC-TCM [45], MLIC++, ELIC [30], and Cheng’20 [16]. All method are quiped with autoregressive context module [52].
+Table 2: Comparisons among different transforms of proposed MLICv2, LIC-TCM liu2023learned, MLIC++, ELIC he2022elic, and Cheng’20 cheng2020learned.
+All method are quiped with autoregressive context module minnen2018image.
+Transform Architecture Cheng’20 cheng2020learned SwinT zhu2022transformerbased ELIC he2022elic MLIC++ LIC-TCM liu2023learned FTIC li2024frequencyaware DCAE lu2025learned MLICv2
+BD-Rate(%) 0.00 -4.61 -5.19 -0.41 -8.91 -8.36 -9.17 -9.44
 Input Image Selected Channel Features of Intra-Slice Local Context and Mean Attention Weights
@@
 The visualized feature maps are normalized to the range [0, 1].
-Context Module MLIC++ MLICv2 Hyperprior Guided Global Context 2D RoPE Context Reweighting Selective Compression
+Table 3: MACs contribution of the introduced modules to the entropy model on the Kodak dataset
+Context Module MLIC++ jiang2023mlicpp MLICv2 HGCP 2D RoPE CR GSC
 GMACs $508.49$ $506.76$ $1.62$ $2.4\times 10^{-5}$ $3.46$ $1.73$
-Table 3: MACs contribution of the introduced modules to the entropy model on the Kodak dataset
-4.4.2 On Forward MACs
-Compared to MLIC+, LIC-TCM [45], our MLICv2 has lower forward MACs.
-Compared to MLIC++, our MLICv2 exhibits slightly higher forward MACs.
+4.4.2 On Forward kMACs/pixel
+Compared to MLIC+, LIC-TCM liu2023learned, our MLICv2 has lower kMACs/pixel.
+Compared to MLIC++, our MLICv2 exhibits slightly higher kMACs/pixel.
 This increase can be attributed to two primary factors: the proposed channel reweighting module and selective compression module.
@@
 4.4.3 On Encoding Time and Decoding Time
-Compared to MLIC and MLIC+ [34], MLICv2 achieves higher encoding and decoding efficiency due to its linear-complexity components.
-Relative to MLIC++[33], it further reduces encoding and decoding time by efficiently skipping many zero elements, which is particularly beneficial for high-resolution images where zeros are more frequent.
+Compared to MLIC and MLIC+ jiang2022mlic, MLICv2 achieves higher encoding and decoding efficiency due to its linear-complexity components.
+Relative to MLIC++jiang2023mlicpp, it further reduces encoding and decoding time by efficiently skipping many zero elements, which is particularly beneficial for high-resolution images where zeros are more frequent.
 Latent refinement, while involving multiple backpropagation steps and increasing encoding time, is optional and suited for scenarios where encoding complexity is less critical.
 Since images are typically encoded once but decoded multiple times, the extra encoding cost is justified by the performance gains.
-Notably, MLICv2+ maintains decoding complexity nearly identical to MLICv2, with both substantially faster than MLIC++, recent WeConvene [23] and DCAE [47].
+Notably, MLICv2+ maintains decoding complexity nearly identical to MLICv2, with both substantially faster than MLIC++, recent WeConvene fu2024weconvene and DCAE lu2025learned.
 4.4.4 On Model Size
@@
 The increase in parameters is minimal, primarily because we deliberately reduce the number of channels for intermediate features in the context modules of MLICv2 to control overall model size.
-Context Modules MLICv2+ Cumulative Benefits Individual Contributions
-Full Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8 Case 9 Case 10 Case 11 Case 12 Case 13 Case 14 Case 15 Case 16
-Latent Refinement
-Selective Compression
-2D RoPE
-Context Reweighting
-Hyperprior Guided Global
-Simple Token Mixing Transform
-Gate
-BD-Rate (%) 0.00 0.17 4.89 5.21 5.77 7.79 8.29 9.58 9.87 4.64 9.52 9.12 7.64 8.77 8.51 9.58 9.87
 Table 4: Complete ablation study on Kodak dataset.
@@
 MLICv2+ is the anchor.
-Cases 1-8 show cumulative benefits, Cases 9-16 show individual contributions.
+Cases 1–7 show cumulative benefits, Cases 9–15 show individual contributions.
+The original Cases 8 and 16 are merged as a separate baseline.
+Context Modules MLICv2+ Cumulative Benefits Individual Contributions Baseline
+Full Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8 Case 9 Case 10 Case 11 Case 12 Case 13 Case 14 Case 15
+ILR
+GSC
+2D RoPE
+CR
+HGCP
+STMT
+Gate
+BD-Rate (%) 0.00 0.17 4.89 5.21 5.77 7.79 8.29 9.58 4.64 9.52 9.12 7.64 8.77 8.51 9.58 9.87
 4.5 Ablation Studies
 4.5.1 Settings
+Figure 13: Distribution of scale values in MLICv2 on Kodak kodak.
 Each ablation cases is optimized for MSE using our established training strategy, with MLICv2+ serving as the anchor.
@@
 The results are presented in Table 4.
-Figure 12: Distribution of scale values in MLICv2 on Kodak [37].
 4.5.2 Analysis of Enhanced Transform
-To demonstrate the superior capacity of our proposed transform, we compare the upper bounds of reconstruction quality achieved by the transforms in MLIC++[33], LIC-TCM[45], and our MLICv2.
+To demonstrate the superior capacity of our proposed transform, we compare the upper bounds of reconstruction quality achieved by the transforms in MLIC++jiang2023mlicpp, LIC-TCMliu2023learned, and our MLICv2.
 The results are shown in Fig.
-10.
+11.
 Our MetaFormer-based design achieves better reconstruction quality than MLIC++ and performs comparably to LIC-TCM, highlighting its stronger representational power despite relying primarily on convolutional operations.
-To demonstrate the effectiveness of the proposed simple token mixing block, we compare the transform module with the transform modules of Cheng’20 [16], SwinT-Charm [87], ELIC [30], and LIC-TCM [45] on rate-distortion performance.
-The entropy model is aligned with spatial autoregressive entropy model [50].
+To demonstrate the effectiveness of the proposed simple token mixing block, we compare the transform module with the transform modules of Cheng’20 cheng2020learned, SwinT-Charm zhu2022transformerbased, ELIC he2022elic, and LIC-TCM liu2023learned on rate-distortion performance.
+The entropy model is aligned with spatial autoregressive entropy model minnen2018joint.
 The results are presented in Table.
@@
 By omitting the encoding of these zero elements, we can significantly accelerate both arithmetic encoding and decoding.
-To demonstrate the effectiveness of proposed guided selective compression on acceleration, the skip ratio, encoding acceleration and decoding acceleration on Kodak [37] and CLIC Pro Val [65] are presented in Table 6.
+To demonstrate the effectiveness of proposed guided selective compression on acceleration, the skip ratio, encoding acceleration and decoding acceleration on Kodak kodak and CLIC Pro Val CLIC2020 are presented in Table 6.
 As image quality improves, the proportion of zero elements that can be skipped decreases, which in turn reduces the acceleration gains for both encoding and decoding.
@@
 To demonstrate the effectiveness of hyperprior-guided global context for the first slice, the attention maps are illustrated in Fig.
-9, which are generally accurate, thereby enhancing the performance.
+10, which are generally accurate, thereby enhancing the performance.
 For slices $\hat{\boldsymbol{y}}^{\geq 1}$, the previous one slice is still employed to predict the global correlations due to the higher bit-rate of each slice compared to the side information.
 To demonstrate the relationship between context reweighting map and image content, we visualize the Mean Attention Weight in Fig. 12.
+Table 5: Model Size Comparison.
 Model
-Xie’21 [75]
-ELIC [30]
-WACNN [88]
-STF [88]
-LIC-TCM [45]
-FTIC [40]
-MLIC+ [34]
-MLIC++ [33]
-WeConvene [23]
-LALIC [20]
-DCAE [47]
+Xie’21 xie2021enhanced
+ELIC he2022elic
+WACNN zou2022the
+STF zou2022the
+LIC-TCM liu2023learned
+FTIC li2024frequencyaware
+MLIC+ jiang2022mlic
+MLIC++ jiang2023mlicpp
+WeConvene fu2024weconvene
+LALIC feng2025linear
+DCAE lu2025learned
 MLICv2
 Parameters (M) 50.0 41.9 75.2 100.0 75.9 69.8 83.4 83.5 105.5 63.2 119.2 84.3
-Table 5: Model Size Comparison.
+Table 6: Skip Ratio, Encoding Acceleration and Decoding Acceleration at different qualities on Kodak and CLIC Pro Val.
 Kodak CLIC Pro Val
@@
 Dec Acceleration (%) 13.61 13.21 12.15 11.97 10.18 9.62 27.52 25.85 25.37 24.21 23.36 20.99
-Table 6: Skip Ratio, Encoding Acceleration and Decoding Acceleration at different qualities on Kodak and CLIC Pro Val.
-Figure 13: Changes in loss, PSNR and bpp during the refinement of latent representation and side information of Kodim01 from Kodak dataset [37].
+Figure 14: Changes in loss, PSNR and bpp during the refinement of latent representation and side information of Kodim01 from Kodak dataset kodak.
 “Initial” denotes the original loss, psnr, bpp when using the latent representation and side information outputed by MLICv2.
@@
 This demonstrates that our context reweighting module effectively emphasizes more informative contexts, supporting its contribution to the performance gains.
-To validate the effectiveness of the proposed 2D RoPE, the 2D RoPE is replaced by original additive relative position embedding [46].
+To validate the effectiveness of the proposed 2D RoPE, the 2D RoPE is replaced by original additive relative position embedding liu2021swin.
 The performance degradation caused by the replacement demonstrates the superiority of the proposed 2D RoPE.
-Inspired by some 2D positional encoding designs [13, 57, 72], which split the feature tensor along the channel dimension and apply positional encoding along the horizontal axis to one half and the vertical axis to the other, we also experimented with a similar strategy.
+Inspired by some 2D positional encoding designs carion2020end, ramachandran2019stand, wang2020axial, which split the feature tensor along the channel dimension and apply positional encoding along the horizontal axis to one half and the vertical axis to the other, we also experimented with a similar strategy.
 Specifically, we divided the input tensor along the channel dimension into two parts and applied RoPE-based rotations along the horizontal axis and vertical axis, respectively.
@@
 In case 1, the latent refinement is removed, which leads to around $4.89\%$ bit-rate increase compared to MLICv2+.
-The increased bit-rate demonstrates the effectiveness of the latent refinement for exploring performance potential.
-The changes of loss, PSNR and bpp are illustrated in Fig.
-13.
-During the refinement, the loss and bpp generally first rise and then fall to lower than the initial loss and bpp, while the PSNR generally keeps rising.
+This significant performance improvement can be attributed to instance-specific adaptation.
+While our MLICv2 model is trained on a diverse dataset to achieve good average performance across various images, each individual image possesses unique characteristics such as texture complexity, edge distribution, and local redundancy patterns.
+ILR enables the latent representation $\hat{\boldsymbol{y}}$ and side information $\hat{\boldsymbol{z}}$ to be specifically optimized for each input image’s rate-distortion characteristics, effectively bridging the gap between generalized encoder mappings ($g_{a}$ and $h_{a}$) and input-specific optimal representations.
+The evolution of loss, PSNR, and bpp during refinement is illustrated in Fig.
+14.
+During the initial refinement steps, both bit-rate and PSNR increase simultaneously, indicating that ILR adjusts latent values near quantization boundaries to better balance rate and distortion for the specific input.
+The subsequent phase exhibits decreasing bit-rate accompanied by further PSNR improvement, demonstrating that the refinement process successfully discovers a more optimal operating point in the rate-distortion space that remains inaccessible to the fixed encoder network.
+This validates the effectiveness of ILR in exploiting the full potential of our architectural framework.
 5 Conclusion
@@
 These innovations collectively establish new state-of-the-art performance benchmarks, with BD-Rate reductions of up to 20% compared to VTM-17.0 Intra while maintaining competitive computational complexity.
-Furthermore, to continue exploring the performance upper bounds via encoding-time scaling, we plan to investigate the transmission of update parameters [67, 59, 49] to the decoder side.
+Furthermore, to continue exploring the performance upper bounds via encoding-time scaling, we plan to investigate the transmission of update parameters tsubota2023universal, shen2023dec, lv2023dynamic to the decoder side.
 Limitations: Although our MLICv2 and MLICv2+ achieve state-of-the-art performance, several limitations warrant future attention.
-First, the performance potential exploration through SGA may lead to bitrate variations [84], requiring practical online encoder optimization strategies for precise bitrate control.
-Second, MLICv2/MLICv2+ may encounter cross-platform decoding issues [7, 29, 61, 79].
-We plan to investigate model quantization techniques [28, 29, 61] and chain coding-based latent compression [79] to address deployment challenges.
+First, the performance potential exploration through SGA may lead to bit-rate variations zhang2024practical, requiring practical online encoder optimization strategies for precise bit-rate control.
+Second, MLICv2/MLICv2+ may encounter cross-platform decoding issues balle2018integer, he2022post, shi2023rate, yang2025learned.
+We plan to investigate model quantization techniques han2015deep, he2022post, shi2023rate and chain coding-based latent compression yang2025learned to address deployment challenges.